### v2

# On Nash resolution of (singular) Lie algebroids

Abstract
Any Lie algebroid $A$ admits a Nash-type blow-up $\mathrm{Nash}(A)$ that sits in a nice short exact sequence of Lie algebroids $0\rightarrow K\rightarrow\mathrm{Nash}(A)\rightarrow\mathcal{D}\rightarrow 0$ with $K$ a Lie algebra bundle and $\mathcal{D}$ a Lie algebroid whose anchor map is injective on an open dense subset. The base variety is a blowup determined by the singular foliation of $A$ considered recently in Omar Mohsen’s [Moh21]. We provide concrete examples. Moreover, we extend the construction to singular subalgebroids in the sense of Androulidakis-Zambon [Zam22].
Keywords: Lie algebroids, singular foliations, singularities.
MSC classes: 53-XX, 53D17, 32S45.
ORCID: $0$-0002-3582-7994
Introduction
We address the following question,
Question 0.1.
Is the slogan “A Lie algebroid splits in a purely manifold part and a purely Lie algebra part” valid ?
There is of course a naive answer: given a Lie algebroid $(A,[\cdot\,,\cdot]_{A},\rho)$ over $M$ with anchor $\rho\colon\Gamma(A)\rightarrow\mathfrak{X}(M)$, there is an exact sequence
[…]
Above, $\ker(\rho)\subseteq\Gamma(A)$ is a totally intransitive object, i.e., a purely Lie algebra object. Let us be more precise. Let $\rho_{m}\colon A_{m}\to T_{m}M$ be the anchor map of $A$ at a point $m\in M$. Consider the collection of Lie algebras $\{\mathrm{Sker}\rho_{m}\}_{m\in M}$ where $u\in\mathrm{Sker}\rho_{m}\subseteq\ker\rho_{m}$ if and only if there exists a local section $\tilde{u}$ of $A$ through $u$ such that $\rho(\tilde{u})=0$. The vector space $\mathrm{Sker}\rho_{m}$ is a Lie ideal of the Lie algebra $\ker\rho_{m}$. Hence, it is a Lie algebra and its Lie bracket {,}$\{\cdot\,,\cdot\}_{\mathrm{Sker}\rho_{m}}$ is given by
$\{a(m),b(m)\}_{\mathrm{Sker}\rho_{m}}:=[a,b]_{A}(m),\qquad a,b\in\ker\rho$
hence depends only on the point $m$. On the base manifold $M$, the image $\mathrm{Im}(\rho)$ is a singular foliation in the sense of [AS09, Cer79, Deb01, LGLS20], so it is entirely “within” the manifold $M$. Hence, it is a purely manifold object. But this answer is not satisfactory, since neither $\ker(\rho)$ nor $\mathrm{Im}(\rho)$ are sections of a vector bundle in general. We will argue that the slogan above is however valid, provided one first applies some canonical blow-up procedure invented by Omar Moshen [Moh21], and in alignment with the conceptual framework of the Nash blow-up of an affine variety[Nob75].
Let us explain the results of this paper. An important class of Lie algebroids are those whose anchor maps are injective on an open dense subset. Those Lie algebroids have been extensively studied by Claire Debord [Deb00, Deb01]. In this paper, we shall call them Debord Lie algebroids. They are naturally the class of Lie algebroids closest to regular foliations, that is, to vector subbundles $F\subseteq TM$ such that $\left[\Gamma(F),\Gamma(F)\right]\subset\Gamma(F)$. In [Deb01], C. Debord proved that Debord Lie algebroids are integrable to smooth groupoids. The image of the anchor map of a Debord Lie algebroid induces a singular foliation, and it is entirely determined by its image through the anchor map. In this sense, it is a purely “manifold object”.
We present a method for desingularizing a Lie algebroid $A$ by finding a pair $(N,p)$ consists of a space $N$ and a map $p\colon N\to M$ such that
  1. 1.
    $p\colon N\to M$ is onto and proper;
  2. 2.
    the restriction $p|_{M_{\mathrm{reg}}}\colon p^{-1}(M_{\mathrm{reg}})\to M_{\mathrm{reg}}$ to regular points $M_{\mathrm{reg}}$ of $A$ is one-to-one
on which the pullback $p^{!}A$ of $A$ admits a Lie algebroid structure $\left(\mathrm{Nash}(A),[\cdot\,,\cdot]_{\mathrm{Nash}(A)},\hat{\rho}\right)$ whose quotient is a Debord algebroid. Also, $\mathrm{Nash}(A)\to A$ is a Lie algebroid morphism. In fact, by construction, $N$ comes equipped with an action of $A$ along $p\colon N\rightarrow M$ in the sense of [Mac05, KSM02, Mok96]. The fundamental assertion presented in Proposition 2.28 of this paper is that we can choose $p\colon N\rightarrow M$ such that there exists a proper Lie algebroid ideal $K\rightarrow N$ of $p^{!}A$ (i.e., with $\Gamma(K)\subseteq\ker(\hat{\rho})$) whose fibers coincide with those of $\ker(\hat{\rho})$ on the open dense subset of regular points. As a consequence, the quotient $\mathrm{Nash}(A)/K$ is a Lie algebroid with injective anchor map on an open dense subset. One of the most important invariant of this construction is the basic singular foliation of $A$ that is, the singular foliation given by the subsheaf of Lie-Rinehart algebras $\mathcal{F}:=\rho_{A}(\Gamma(A))\subseteq\mathfrak{X}(M)$. In other words, $N$ does not depend on the Lie algebroid $A$ but only on its basic singular foliation. Importantly, this result remains valid even when the underlying structure $(A\rightarrow M,[\cdot\,,\cdot]_{A},\rho_{A})$ is merely an almost Lie algebroid, i.e., without the requirement that $[\cdot\,,\cdot]_{A}$ adheres to the Jacobi identity: the anchor map still has to be a morphism of brackets.
In some cases, our blow-up space $N$ is precisely the usual blow-up along the singular part of the singular foliation $\mathcal{F}=\rho_{A}(\Gamma(A))$ induced by the Lie algebroid. We must admit that very often our $N$ is not a smooth manifold. However, our Lie algebroids structures on $N$ admits leaves that are smooth manifolds. But if the closed subset of singular points $M_{\mathrm{sing}}$ is a smooth manifold and $\ker\rho$ is finitely generated, then $N$ is a smooth manifold [Ser89].
The construction of $N$ relies on a very classical method due to the mathematician J. Nash [Nob75], mainly used in algebraic geometry for desingularisation. The idea of applying this method to singular foliations is not totally new. As far as I know, A. Sertöz Sinan was the first who adapted it on foliations in his PhD [Ser89] in order to compute the Baum-Bott residues of singular holomorphic foliations, see also [PBTS00]. In fact, A. Sertöz Sinan did more, he applied Nash in the realm of coherent subsheaves of locally free sheaves and generalized the works of Nobile [Nob75]. The idea of Nash blow-up $\mathrm{Nbl}_{\mathcal{F}}(M)$ of a singular foliated manifold $(M,\mathcal{F})$ in the sense of [AS09] was used by O. Mohsen [Moh21, Moh22] and consists in replacing every singular point by the limiting positions of the tangent spaces of the leaves in neighborhoods of regular points. In this paper we do not apply the Nash method directly on the tangent space of our singular foliation $\mathcal{F}\subseteq\mathfrak{X}(M)$ but on the kernel of the anchor map $\rho\colon A\rightarrow TM$. The choice of considering the kernel allows generalizing the Nash construction to Lie $\infty$-algebroid that are universal in the sense of [LGLS20, LGL22]. This is discussed by the author in [Lou23]. The latter provides a class of applications of this method on solvable singular foliations, the author described a sequence of Nash-blow-ups associated to a geometric resolution of a singular foliation in the sense of [LGLS20].
These constructions enable us, in passing, to efficiently compute several instances of blowup of a singular foliation.
Also, we show in Proposition 3.8 that the space of all limits of the tangent spaces of the regular symplectic leaves of a Poisson manifold is isomorphic to the space of all the limits of the kernel of the associated contangent algebroid on regular points. This justifies the name “Nash blow-up”. In the Poisson case, we give an example where the Poisson bivector field degenerates under the blow-up procedure, see Example 3.9. Although, all its Hamiltonian vector fields remain well-defined after this procedure.
Moreover, we claim that our method can be applied to many “Lie algebroids” type structures, such as “almost Lie algebroids” [PP19], “left-symmetric algebroids” in the sense of [LSBC16] or “Hom-Lie algebroids” [LGT13, ZYW17] and so on. And one could easily replace with no additional assumptions the (Lie) bracket of our main theorem with the corresponding operation.
The structure of the paper is the following. In Section 1, we recall some basics on singular subalgebroids and singular foliations. In Section 2, we introduce the Nash construction and state the main results of the paper with their proofs. In Section 3, we end the paper with some examples from singular foliations and Poisson geometry.
Acknowledgements. This paper is part of a post-doc fellowship at Jilin University and Göttingen University. I acknowledge important discussions with Jun Jiang and Camille Laurent-Gengoux at early stages of the project. Finally, I would like to thank Yunhe Sheng for his technical assistance at the Mathematics Laboratory of Jilin University in the realization of this paper.
1 Definitions
Throughout the paper, $M$ stands for a smooth/analytic/complex manifold or an algebraic variety over $\mathbb{K}\in\{\mathbb{R},\mathbb{C}\}$. We denote by $\mathcal{O}_{M}$ the algebra of smooth/holomorphic or polynomial functions on $M$ depending on the context. Also, $\mathbb{K}$ stands for $\mathbb{R}$ or $\mathbb{C}$.
Let $(E,\left[\cdot\,,\cdot\right]_{E},\rho_{E})$ be a Lie algebroid over a manifold $M$.
1.1 Singular subalgebroids
The following concept was first introduced in [Zam22].
Definition 1.1.
A singular subalgebroid of $E$ is a $\mathcal{O}_{M}$-submodule $\mathcal{F}\subseteq\Gamma(E)$ which is
  1. 1.
    involutive, i.e., $[\mathcal{F},\mathcal{F}]_{E}\subseteq\mathcal{F}$;
  2. 2.
    locally finitely generated [^1] In the complex or real analytic case this is automatically satisfied [Tou68]., i.e., every point $m\in M$ admits an open neighborhood $\mathcal{U}$ on which there exists $\xi_{1},\dots,\xi_{k}\in\mathcal{F}|_{\mathcal{U}}\subseteq\Gamma(E)|_{\mathcal{U}}$ such that for every $\mathcal{V}\subset\mathcal{U}$, the restrictions of $\xi_{1},\dots,\xi_{k}$ to $\mathcal{V}$ generate $\mathcal{F}_{\mathcal{V}}$ as a $\mathcal{O}_{\mathcal{V}}$-module.
Remark 1.2.
Notice that Definition 1.1 means that a singular subalgebroid $\mathcal{F}\subseteq\Gamma(E)$ is a locally finitely generated Lie-Rinehart algebra [Hue03] whose Lie bracket and anchor map are the restriction of the Lie bracket $[\cdot\,,\cdot]_{E}$ and the anchor map $\rho_{E}$ to $\mathcal{F}$.
Example 1.3 (Singular foliations).
One of an important class of singular subalgebroids for this paper is the class of singular foliations in the sense of [Her62, AS09, Deb01, LGLR24]. A singular foliation $\mathcal{F}$ on $M$ is a singular subalgebroid of $E=TM$. Here are some important features of singular foliations in the smooth/complex cases.
    Singular foliation admits leaves [Her62]: there exists a partition of $M$ into immersed submanifolds (possibly of non-constant dimension) called leaves such that for all $m\in M$, the image of the evaluation map $\mathcal{F}\to T_{m}M$ is the tangent space of the leaf through $m$.
    Singular foliations are self-preserving: the flow $\phi^{X}_{t}$ of vector fields $X\in\mathcal{F}$, whenever defined, preserves $\mathcal{F}$ [Her62, AS09, GY18], i.e., $\forall\,m\in M,\exists\,\varepsilon>0$ and a neighborhood $\mathcal{U}$ of $m$ such that $\forall t\in]-\varepsilon,\varepsilon[$
    ${(\phi^{X}_{t}|_{\mathcal{U}})}_{*}(\mathcal{F})=\mathcal{F}_{\phi_{t}^{X}(\mathcal{U})}.$
Example 1.4 (Debord singular subalgebroids).
Here is a special class of singular subalgebroids. A singular subalgebroid $\mathcal{F}$ of $E$ is said to be Debord or projective if there is a vector bundle morphism $A\stackrel{{\scriptstyle\rho}}{{\rightarrow}}E$ covering the identity of $M$ such that $\rho(\Gamma(A))=\mathcal{F}$ and which is injective on an open dense subset. In that case, there is a Lie bracket on $A$ such that $\rho$ is a Lie algebroid morphism. Therefore, $\mathcal{F}$ arise as the image of a Lie algebroid.
For more examples, see [Zam22].
1.1.1 $E$-valued almost Lie algebroids and singular subalgebroids
Most of our method will be valid for almost Lie algebroids. Let us recall some definitions.
Definition 1.5 ([PP19, LGLS20]).
Let $\mathcal{F}\subseteq\Gamma(E)$ be a singular subalgebroid. An $E$-valued almost Lie algebroid $(A,\left[\cdot\,,\cdot\right]_{A},\rho_{A})$ over $\mathcal{F}$ is a vector bundle morphism $A\stackrel{{\scriptstyle\rho_{A}}}{{\rightarrow}}E$ (referred as the anchor map) such that $\rho_{A}(\Gamma(A))=\mathcal{F}$; together with a skew-symmetric bilinear map $\left[\cdot\,,\cdot\right]_{A}\colon\Gamma(A)\wedge\Gamma(A)\rightarrow\Gamma(A)$ (referred to as the bracket) satisfying for all $a,b\in\Gamma(A)$ and $f\in\mathcal{O}_{M}$
  1. 1.
    $\rho_{A}(\left[a,b\right]_{A})=[\rho_{A}(a),\rho_{A}(b)]_{E}$;
  2. 2.
    $\left[a,fb\right]_{A}=\rho_{A}(a)[f]b+f\left[a,b\right]_{A}$.
$\mathcal{F}$ shall be called the basic singular subalgebroid of $(A,\left[\cdot\,,\cdot\right]_{A},\rho_{A})$. If $E=TM$ we simply say “almost Lie algebroid”.
If $\left[\cdot\,,\cdot\right]_{A}$ satisfies the Jacobi identity, i.e., for $a,b,c\in\Gamma(A)$
$[[a,b]_{A},c]_{A}+[[c,a]_{A},b]_{A}+[[b,c]_{A},a]_{A}=0,$ (1)
then $A$ is a Lie algebroid over $M$ with anchor map $\rho_{E}\circ\rho_{A}$.
Remark 1.6.
Every singular subalgebroid is the image of an $E$-valued almost Lie algebroid of on any open subset whenever it is finitely generated. For instance, a singular subalgebroid is finitely generated if and only is it is the image of some $E$-valued almost Lie algebroid, see [LGLS20], Proposition 2.1.13, p.99.
Proposition 1.7.
For any two almost algebroids $(A_{1},\left[\cdot\,,\cdot\right]_{A_{1}},\rho_{A_{1}})$ and $(A_{2},\left[\cdot\,,\cdot\right]_{A_{2}},\rho_{A_{2}})$ over a singular subalgebroid $\mathcal{F}\subseteq\Gamma(E)$, there exist morphisms of vector bundles
${A_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${A_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$ (2)
such that
  1. 1.
    $\rho_{A_{2}}\circ\Phi=\rho_{A_{1}}$ and $\rho_{A_{1}}\circ\Psi=\rho_{A_{2}}$.
  2. 2.
    $\rho_{A_{2}}(\Phi\circ\Psi-\mathrm{id})=0$ and $\rho_{A_{1}}(\Psi\circ\Phi-\mathrm{id})=0$.
In particular, $[\Phi(a),\Phi(b)]_{A_{2}}-\Phi([a,b]_{A_{1}})\in\ker\rho_{A_{2}}$ for all $a,b\in\Gamma(A_{1})$. Also, we have $[\Psi(a),\Psi(b)]_{A_{1}}-\Psi([a,b]_{A_{2}})\in\ker\rho_{A_{1}}$ for all $a,b\in\Gamma(A_{2})$.
Proof.
For any pair of $E$-valued almost Lie algebroids $(A,[\cdot\,,\cdot]_{A_{1}},\rho_{A_{1}})$ and $({A_{2}},[\cdot\,,\cdot]_{A_{2}},\rho_{A_{2}})$ over $\mathcal{F}$, there exists (by projectivity of the module of sections of vector bundles) vector bundle morphisms
${(A,[\cdot\,,\cdot]_{A},\rho_{A})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${(A_{2},[\cdot\,,\cdot]_{A_{2}},\rho_{A_{2}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$
that satisfies item 1. The other properties are automatically verified. ∎
Let us give a crucial result on linear vector fields on $E$-valued almost Lie algebroids. These techniques are used in [BLM19]. Recall that for $q\colon A\rightarrow M$ be a vector bundle over $M$ a linear vector field on $A$ is a couple $(\xi,X)\in\mathfrak{X}(A)\times\mathfrak{X}(M)$ such that the following diagram
${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\xi}$$\scriptstyle{q}$${TA\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{Tq}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{X}$${TM}$
commutes. Equivalently,
$\xi[\Gamma(A^{*})]\subseteq\Gamma(A^{*})\;\;\text{and}\;\;\xi[q^{*}\mathcal{O}_{M}]\subseteq q^{*}\mathcal{O}_{M}.$ (3)
For more details and properties on linear vector fields on vector bundles, see [Mac05].
Proposition 1.8.
Let $(A,\left[\cdot\,,\cdot\right]_{A},\rho_{A})$ be an $E$-valued almost Lie algebroid over a singular Lie algebroid $\mathcal{F}\subseteq\Gamma(E)$. Let $e\in\mathcal{F}=\rho(\Gamma(A))$ and $X=\rho_{E}(e)$, there exists a linear vector field $\xi_{e}$ on $q\colon A\rightarrow M$ and a linear vector field $\nu_{X}$ on $q^{\prime}\colon E\to M$ such that
  1. 1.
    $\xi_{e}$ is $q$-related to $X$ and $\nu_{X}$ is $q^{\prime}$-related to $X$,
  2. 2.
    their flows at time $t$ are defined if and only if the flow of $X$ at time $t$ is defined,
  3. 3.
    the flows $\phi^{\xi_{e}}_{t},\;\phi^{\nu_{X}}_{t}$ whenever defined are isomorphism of vector bundles that make the following diagram commutes,
[…]
Proof.
Let $e\in\mathcal{F}$ and $X=\rho_{E}(e)$. Choose $a\in\Gamma(A)$ a section of $q\colon A\to M$ such that $\rho(a)=e$. The Lie derivatives $\mathcal{L}_{a},\,\mathcal{L}_{e}$ defined by
  1. 1.
    $\mathcal{L}_{a}[q^{*}f]:$ $=q^{*}(X[f]),\;\forall\;f\in\mathcal{O}_{M},$
    $\langle\mathcal{L}_{a}[\alpha],b\rangle$ $=X[\langle\alpha,b\rangle]-\langle\alpha,[a,b]_{A}\rangle,\quad\forall\alpha\in\Gamma(A^{*}),\;b\in\Gamma(A).$
  2. 2.
    $\mathcal{L}_{e}[q^{\prime*}f]:$ $=q^{\prime*}(X[f]),\;\forall\;f\in\mathcal{O}_{M},$
    $\langle\mathcal{L}_{e}[\beta],e^{\prime}\rangle$ $=X[\langle\beta,e^{\prime}\rangle]-\langle\beta,[e,e^{\prime}]_{E}\rangle,\quad\forall\beta\in\Gamma(E^{*}),\;e^{\prime}\in\Gamma(E).$
define linear vector fields $\xi_{e}\in\mathfrak{X}(A),\;\nu_{X}\in\mathfrak{X}(E)$ which are respectively $q$-related to $X$ and $q^{\prime}$-related to $X$. Notice that these linear vector fields depend respectively on the brackets $[\cdot\,,\cdot]_{A},\;[\cdot\,,\cdot]_{E}$ and the vector field $X$. This proves item 1. An argument to prove item 2 and 3 is the following: the vector field $(\xi_{e}+\nu_{X},X)$ is a linear vector field on $A\oplus E$ which is tangent to the graph of $\rho_{A}\colon A\to E$:
$\left\{(u,\rho_{A}(u))|u\in A\right\}\subset A\oplus E.$
Hence, the diagram (4) commutes. For a complete proof of item 2 and 3 see [LGLR24], Lemma 2.1.19. p. 61. ∎
Remark 1.9.
When $E=TM$, we have $\phi_{t}^{\nu_{X}}=T\phi_{t}^{X}$ in Diagram (4).
1.1.2 Isotropy Lie algebras of a singular subalgebroid
We start with the following definition
Definition 1.10 (Strong-kernel).
Let $\eta\colon A\rightarrow E$ be a vector bundle morphism covering the identity map. The strong-kernel $\mathrm{Sker}\,\eta_{m}$ of $\eta$ at $m\in M$ is the subvector space of $\ker\eta_{m}$ defined as $u\in A_{m}$ such that there is an open neighborhood $\mathcal{U}$ of $m$ and a local section $\tilde{u}\in\Gamma_{\mathcal{U}}(A)$ through $u$ so that $\eta(\tilde{u})=0$.
Here are some easy properties of the strong-kernel
Proposition 1.11.
[LGLR24, AS09] Let $\eta\colon A\rightarrow E$ be a vector bundle morphism covering the identity map.
  1. 1.
    The map $m\mapsto\dim(\mathrm{Sker}\,\eta_{m})$ is lower semi-continuous and $m\mapsto\dim(\ker\eta_{m})$ is upper semi-continuous.
  2. 2.
    $\mathrm{Sker}\,\eta_{m}=\ker\eta_{m}$ if and only if there is open neighborhood $\mathcal{U}$ of $m$ such that $\mathrm{Sker}\,\eta_{m^{\prime}}=\ker\eta_{m^{\prime}}$ for all $m^{\prime}\in\mathcal{U}$.
  3. 3.
    The open dense subset where $m\mapsto\ker\eta_{m}$ and $m\mapsto\mathrm{Im}\eta_{m}$ are locally constant is the set of points $m\in M$ such that $\mathrm{Sker}\,\eta_{m}=\ker\eta_{m}$.
The notion that we now introduce generalizes the original notion of isotropy Lie algebras for singular foliations in [AS09].
Let $\mathcal{F}\subseteq\Gamma(E)$ be a singular subalgebroid. We may assume the $\mathcal{F}$ is finitely generated, so that it is the image of an $E$-valued almost Lie algebroid $(A,\left[\cdot\,,\cdot\right]_{A},\rho)$. For $m\in M$, the skew-symmetric bracket $\left[\cdot\,,\cdot\right]_{A}$ induces a bilinear map
$\left[\cdot\,,\cdot\right]_{A,m}\colon\wedge^{2}\ker\rho_{m}\rightarrow\ker\rho_{m}$
defined for all $u,v\in\ker\rho_{m}$ by $\left[u\,,v\right]_{A,m}:=\left[\widetilde{u}\,,\widetilde{v}\right]_{A}$ where $\widetilde{u},\widetilde{v}$ are respectively local sections through $u$ and $v$ respectively in an open neighborhood of $m$. It is easily checked that this pointwise bracket is well-defined, that is $\left[\cdot\,,\cdot\right]_{A}$ depends on the $1$-jet at $m$ of the sections of $A$. Furthermore, the strong kernel $\mathrm{Sker}\,\rho_{m}$ at $m$ satisfies $[{\mathrm{Sker}}\,\rho_{m},{\ker}\rho_{m}]_{A,m}\subset{\mathrm{Sker}}\,\rho_{m}$. This implies that the skew-symmetric bilinear map $[\cdot,\cdot]_{A,m}$ goes to the quotient to a bilinear map,
{,}$\{\cdot,\cdot\}_{m}\colon\wedge^{2}\frac{\ker\rho_{m}}{\mathrm{Sker}\,\rho_{m}}\longrightarrow\frac{\ker\rho_{m}}{\mathrm{Sker}\,\rho_{m}}$ (5)
which turns out to be a Lie bracket [LGLR24].
Definition 1.12.
The isotropy Lie algebra of $\mathcal{F}$ at $m$ is the quotient
$\mathfrak{g}_{m}(\mathcal{F})=\frac{\ker\rho_{m}}{\mathrm{Sker}\,\rho_{m}}$
together with the Lie bracket {,}$\{\cdot\,,\cdot\}_{m}$.
The following proposition is a direct consequence of Proposition 1.7
Proposition 1.13.
The isotropy Lie algebra $\mathfrak{g}_{m}(\mathcal{F})$ of a singular subalgebroid $\mathcal{F}$ at $m$ is independent of the choice of an $E$-valued almost Lie algebroid over $\mathcal{F}$.
Proof.
This is a direct consequence of Proposition 1.7. ∎
Remark 1.14 (Androulidakis-Skandalis isotropy Lie algebras).
Let $\mathcal{\mathcal{F}}\subseteq\Gamma(E)$ be a singular subalgebroid and $m\in M$. Denote by $\mathcal{I}_{m}$ be the ideal of vanishing functions at $m$ and $\mathcal{F}(m)$ the Lie algebra of sections of $\mathcal{F}$ vanishing at $m$ which contains the ideal $\mathcal{I}_{m}\mathcal{F}$. Assume that $\mathcal{F}$ is the basic singular subalgebroid of some $E$-valued almost Lie algebroid $(A,[\cdot\;,\cdot]_{A},\rho)$. There is a natural Lie algebra isomorphism $\mathfrak{g}_{m}(\mathcal{F})\simeq\frac{\mathcal{F}(m)}{\mathcal{I}_{m}\mathcal{F}}$, see [LGLR24]. When $\mathcal{F}\subseteq\mathfrak{X}(M)$ is a singular foliation, $\mathfrak{g}_{m}(\mathcal{F})$ is the isotropy Lie algebra of $\mathcal{F}$ at $m$ in the sense of [AS09].
2 Nash-blow-up of an $E$-valued almost Lie algebroid
2.1 Grassmann bundle
Convention 2.1.
For our current purpose, we consider the Grassmannian of all sub-spaces of co-dimension $k\in\mathbb{N}$ in $\mathbb{C}^{N}$ rather than the set of subspaces of a given dimension. It is convenient to denote this manifold by $\mathrm{Grass}_{-k}(\mathbb{C}^{N})$ (notice the use of a minus sign). In other words, we set $\mathrm{Grass}_{-k}(\mathbb{C}^{N}):=\mathrm{Grass}_{N-k}(\mathbb{C}^{N})$.
Now, we consider this construction pointwise on a given vector bundle over a manifold. The next definition makes sense in the smooth or complex cases without adaptation.
Definition 2.2.
Let $A\rightarrow M$ be a vector bundle of rank $d$ over a manifold $M$. Let $k\leq\mathrm{rk}(A)=d$. The disjoint union:
$\mathrm{Grass}_{-k}({A}):=\coprod_{x\in M}\mathrm{Grass}_{-k}(A_{x})$
comes equipped with a natural manifold structure. Also,
$P\colon\mathrm{Grass}_{-k}(A)\longrightarrow M$ (6)
is a fibration with typical fiber $\mathrm{Grass}_{-k}(\mathbb{K}^{d})$. It is called $-k$-th Grassmann bundle of $A$.
The following straightforward statement will be used in the sections. Again, it is valid in both smooth and complex cases.
Proposition 2.3.
Let $A\longrightarrow M$ be a vector bundle over $M$. A linear vector field $(\xi,X)$ on $A$ with base vector field $X\in\mathfrak{X}(M)$ induces a vector field on the Grassmann bundle $P\colon\mathrm{Grass}_{-k}(A)\longrightarrow M$ which is $P$-related to $X$.
The tautological bundle. There is a natural and important exact sequence of vector bundles over $\mathrm{Gr}_{-k}(A)$, for $A\to M$ a vector bundle and $k$ an integer $\leq{\mathrm{rk}}(A)$.
  1. 1.
    Let $P^{!}A$ be the pull-back of the vector bundle $A\to M$ to $\mathrm{Grass}_{-k}(A)$.
  2. 2.
    The tautological subbundle $\tau_{A}^{-k}\longrightarrow\mathrm{Grass}_{-k}(A)$ is the vector subbundle of $P^{!}A$ whose fiber at a point $V_{x}\in\mathrm{Grass}_{-k}(A)$ is precisely the vector space $V_{x}$, seen a subspace of $\left(P^{!}A\right)_{V_{x}}=A_{x}$ (the $-k$ being a reminder that it is of corank $k$ in $P^{!}A$).
  3. 3.
    We denote by $Q_{A}^{k}$ and call the tautological quotient bundle the bundle over $\mathrm{Grass}_{-k}(A)$ obtained by taking the quotient of the first bundle by the second one, i.e., $Q_{A}^{k}:=P^{!}A/\tau_{A}^{-k}$ (the $k$ being a reminder that this quotient has rank $k$).
These three vector bundles fit into the short exact sequence
[…]
2.2 The blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$: Nash construction
Omar Mohsen [Moh21] developed a blow-up method for an action Lie algebroid associated with a Lie algebra $\mathfrak{g}$ acting on a manifold $M$. Additionally, he employed a comparable approach to define the blow-up space of any arbitrary singular foliation, without using additional underlying structures. In this section, we exploit his concept to build a blow-up space using anchored bundles. This facilitates the application of this methodology to any object resembling a Lie algebroid. As said in the introduction, the construction is valid for arbitrary vector bundle morphisms, which is a particular case of A. Sertöz Sinan [Ser89] applied in the realm of coherent subsheaves of locally free sheave. This point of view allows us to apply our result to a larger class than the one of singular foliations, namely, singular subalgebroids in the sense of Androulidakis-Zambon [Zam22].
Let $A,E$ be vector bundles over $M$. Consider a morphism of vector bundles
${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\rho}$${E\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${M}$
Denote by $\mathcal{F}:=\rho(\Gamma(A))\subseteq\Gamma(E)$ the subsheaf of $\mathcal{O}_{M}$-submodule (after sheafification).
Convention 2.4.
We shall denote by $M_{\mathrm{reg}}\subseteq M$ the open dense subset where $x\mapsto\ker\rho_{x}$ and $x\mapsto\mathrm{Im}\,\rho_{x}$ are locally constant. Points of $M_{\mathrm{reg}}$ shall be called regular and points in $M\setminus M_{\mathrm{reg}}$ are called singular. In the complex case, the dimension of $\mathrm{Im}\,\rho_{x}\subseteq E_{x}$ is the same for all points $x\in M_{\mathrm{reg}}$. In the smooth case, we will from now assume that $\rho$ is of constant rank on $M_{\mathrm{reg}}$. This happens in particular when $\mathcal{F}$ admits a geometric resolution of finite length [LGLS20]. Let $r$ be the common dimension of $\mathrm{Im}\,\rho_{x}\subseteq E_{x}$ on $M_{\mathrm{reg}}$. The restriction $\mathcal{F}|_{M_{\mathrm{reg}}}\subseteq E|_{M_{\mathrm{reg}}}$ is a vector subbundle of rank $r$, called the regular part of $\mathcal{F}$.
Remark 2.5.
When $E=TM$ and $\mathcal{F}\subseteq\mathfrak{X}(M)$ is a singular foliation, $M_{\mathrm{reg}}$ is the union of the regular leaves of $\mathcal{F}$. More generally, when $E$ is a Lie algebroid and $\mathcal{F}\subseteq\Gamma(E)$ is a singular subalgebroid, $M_{\mathrm{reg}}$ is the set of points $m\in M$ whose isotropy Lie algebra vanishes, i.e., such that $\mathfrak{g}_{m}(\mathcal{F})=0$. Moreover, the restriction $\mathcal{F}|_{M_{\mathrm{reg}}}\subseteq E|_{M_{\mathrm{reg}}}$ is a Lie subalgebroid of $E|_{M_{\mathrm{reg}}}$.
Construction of the blow-up space. Notice that for every $x\in M_{\mathrm{reg}}$
  1. 1.
    $\mathrm{Im}\,\rho_{x}$ is a point of the Grassmannian $\mathrm{Grass}_{r}(E_{x})$, and
  2. 2.
    $\ker\rho_{x}$ is a point of $\mathrm{Grass}_{-r}(A_{x})$.
Consider now the Grassmann bundle $\mathrm{Grass}_{r}(E)$, and the Grassmann bundle $\mathrm{Grass}_{-r}(A)$. We denote by $P$, in both cases, their natural projections onto $M$. Consider the two natural sections of $P$ defined on $M_{\mathrm{reg}}$ by:
  1. 1.
    $\sigma_{\mathrm{Im}}\colon M_{\mathrm{reg}}\longrightarrow\mathrm{Grass}_{r}(E),\,x\longmapsto\mathrm{Im}\,\rho_{x}$, and
  2. 2.
    $\sigma_{\ker}\colon M_{\mathrm{reg}}\longrightarrow\mathrm{Grass}_{-r}(A),\,x\longmapsto\ker\rho_{x}$.
Then we define
  1. 1.
    $\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}}(M):=\overline{\sigma_{\mathrm{Im}}(M_{\mathrm{reg}})}$ to be the closure of the image of the section $\sigma_{\mathrm{Im}}$ in $\mathrm{Grass}_{r}(E)$, and
  2. 2.
    $\mathrm{Nbl}^{\ker}_{\mathcal{F}}(M):=\overline{\sigma_{\ker}(M_{\mathrm{reg}})}$ to be the closure of the image of the section $\sigma_{\ker}$ in $\mathrm{Grass}_{-r}(A)$.
We shall denote by $p$ the restriction of $P$ to both $\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}}(M)$ and $\mathrm{Nbl}^{\ker}_{\mathcal{F}}(M)$.
Remark 2.6.
Notice that in [Lou23], a sequence, depending on $n\in\mathbb{N}$, of “blow-ups” is constructed, which for $n=0$ gives back $\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}}(M)$, and for $n=1$ gives $\mathrm{Nbl}^{\ker}_{\mathcal{F}}(M)$.
In section 3.2, we have a special class of examples where $\mathrm{Nbl}^{\ker}_{\mathcal{F}}(M)\simeq\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}}(M)$, namely Poisson structures.
Convention 2.7.
From now on, we will only consider $\mathrm{Nbl}^{\ker}_{\mathcal{F}}(M)$, that we simply denote it by $\mathrm{Nbl}_{\mathcal{F}}(M)$ and shall be called the Nash blow-up of $M$ along $\mathcal{F}$. This notation, which only mentions $(M,\mathcal{F})$, is justified in Proposition 2.15 below.
Remark 2.8.
Notice that
  1. 1.
    Intuitively, for $x\in M$, $p^{-1}(x)=\mathrm{Nbl}_{\mathcal{F}}(M)\cap P^{-1}(x)$ is the set of all possible limits of the subspaces $\ker\rho_{y}$ when $y\in M_{\mathrm{reg}}$ converges to $x$. More precisely, for any $x\in M$, there is a an open neighborhood $\mathcal{U}\subset M$ of $x$ such that $\mathrm{Grass}_{-r}(A)\simeq\mathcal{U}\times\mathrm{Grass}_{-r}(\mathbb{K}^{\mathrm{rk}(A)})$. By construction,
    $p^{-1}(x)=\left\{V\subset A_{{}_{x}}\;\middle|\;\exists\,(x_{n})\in M_{\mathrm{reg}}^{\mathbb{N}},\,\text{such that},\,\;\ker\rho_{x_{n}}\underset{n\to+\infty}{\longrightarrow}V\text{ as }x_{n}\underset{n\to+\infty}{\longrightarrow}x\right\}.$
  2. 2.
    For any open subset $\mathcal{U}$ of $M$, $\sigma_{\ker}(\mathcal{U}\cap M_{\mathrm{reg}})$ is dense in $p^{-1}(\mathcal{U})$.
We gather in the following proposition some easy but important features of that construction.
Proposition 2.9.
Let $\rho\colon A\to E$ be a vector bundle morphism. In the smooth case, we assume that all regular leaves have the same dimension $r$. The map $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$ (referred as the blow-up map) satisfies the following properties:
  1. 1.
    $p$ is proper and surjective. In particular, for each point $x\in M$, the fiber $p^{-1}(x)$ is a non-empty compact set.
  2. 2.
    For every $x\in M$ and $V\in p^{-1}(x)$, $\mathrm{Sker}\,\rho_{x}\subseteq V\subseteq\ker\rho_{x}$.
  3. 3.
    For every $x\in M$ and $V\in p^{-1}(x)$, the image $[V]$ of $V$ in the quotient vector space
    $\mathfrak{g}_{x}(\mathcal{F})=\frac{\ker\rho_{x}}{\mathrm{Sker}\,\rho_{x}}$
    is a vector space of codimension $r-\dim(\mathrm{Im}\,\rho_{x})$. Also, $V\to[V]$ is an injective map from $p^{-1}(x)$ to ${\mathrm{Grass}}_{-(r-\dim(\mathrm{Im}\,\rho_{x}))}(\mathfrak{g}_{x}(\mathcal{F}))$.
  4. 4.
    For every $x\in M_{\mathrm{reg}}$, $p^{-1}(x)=\{\ker\rho_{x}\}$ is reduced to a single point in $\mathrm{Grass}_{-r}(A)$. Also, $p^{-1}(M_{\mathrm{reg}})$ is a smooth manifold and the restriction $p\colon p^{-1}(M_{\mathrm{reg}})\longrightarrow M_{\mathrm{reg}}$ is invertible [^2] Invertible here means: diffeomorphism, in the smooth case, bi-holomorphism, in the complex case..
Proof.
The proof follows the same lines as Proposition 2.5 in [Lou23] for $n=1$. ∎
Remark 2.10.
Despite the extremely pleasant properties, the blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ has a major problem: it is not a sub-manifold in general. However, since $\mathrm{Nbl}_{\mathcal{F}}(M)\subset\mathrm{Grass}_{-r}(A)$ is defined as a closed subset, it inherits a “good” algebra of smooth functions and can be seen as a nice differential space, see Section 2.2.1. In addition, we show that it admits in the case we are interested in a partition into smooth manifolds which form in fact the leaves of the pull-back foliation [^3] We make sense of singular foliation on $\mathrm{Nbl}_{\mathcal{F}}(M)$ in Section 2.2.1.of $\mathcal{F}$ see Section 2.3.
We have the following smoothness result, which is a direct consequence of [Ser89], Corollary 1.2.
Proposition 2.11.
If $\ker\rho\subset\Gamma(A)$ is a finitely generated $\mathcal{O}_{M}$-submodule and the singular locus $M\setminus M_{\mathrm{reg}}$ is a smooth submanifold of $M$, then $\mathrm{Nbl}_{\mathcal{F}}(M)$ is the usual blow-up of $M$ along $M\setminus M_{\mathrm{reg}}$. Hence, $\mathrm{Nbl}_{\mathcal{F}}(M)$ is a smooth manifold.
Remark 2.12.
The first assumption of Proposition 2.11 is fulfilled in particular if $\mathcal{F}$ admits a geometric resolution in the sense of [LGLS20]. For a more general statement than Proposition 2.11 see [Lou23].
Until now, we had defined the Nash blow-up for any vector bundle morphism $\rho\colon A\to E$. Now let us add an $E$-valued almost Lie algebroid bracket on $A$ so that the image $\mathcal{F}=\rho(\Gamma(A))$ be a singular subalgebroid of $E$.
The proof of the following proposition follows the same lines of Proposition 2.5 in [Lou23].
Proposition 2.13.
Let $\rho\colon A\to E$ be a vector bundle morphism. We assume that $(A\rightarrow M,\rho)$ is equipped with a skew-symmetric bracket that makes it an $E$-valued almost Lie algebroid with anchor map $\rho$. In the smooth case, we assume that all regular leaves have the same dimension $r$.
  1. 1.
    For every $x\in M$ and $V\in p^{-1}(x)$, $V$ is stable under the skew-symmetric bracket of $\ker\rho_{m}$.
  2. 2.
    For every $x\in M$ and $V\in p^{-1}(x)$, the image $[V]$ of $V$ in the isotropy Lie algebra $\mathfrak{g}_{x}(\mathcal{F})$ at $x$ is a Lie subalgebra of codimension $r-\dim(\mathrm{Im}\,\rho_{x})$. Also, $V\to[V]$ is an injective map from $p^{-1}(x)$ to ${\mathrm{Grass}}_{-(r-\dim(\mathrm{Im}\,\rho_{x}))}(\mathfrak{g}_{x}(\mathcal{F}))$.
Remark 2.14.
Let $(A,[\cdot\,,\cdot]_{A},\rho)$ be an $E$-valued almost Lie algebroid. The Jacobiator $J\colon\wedge^{3}A\rightarrow A$ (the failure of having Jacobi identity) fits in non-exact complex
$\wedge^{3}A\stackrel{{\scriptstyle J}}{{\rightarrow}}A\stackrel{{\scriptstyle\rho}}{{\rightarrow}}E$ (8)
Here is a natural question which can be the object of another study: what kind of structure we obtain if we apply the Nash construction to the vector bundle morphism $J$?
Now we justify the notation $\mathrm{Nbl}_{\mathcal{F}}(M)$: it does not mention the anchor map $\rho\colon A\rightarrow E$, because, indeed, it does not depend on the $E$-valued almost Lie algebroid structure. More precisely,
Proposition 2.15.
$\mathrm{Nbl}_{\mathcal{F}}(M)$ does not depend on the choice of an $E$-valued almost Lie algebroid over $\mathcal{F}$, i.e., the sets associated to two different anchored bundles are canonically isomorphic.
Proof.
Let $\mathcal{F}\subseteq\Gamma(E)$ a singular subalgebroid and $A,A^{\prime}$ be two $E$-valued almost Lie algebroids over $\mathcal{F}$. Let
${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Phi}$${A^{\prime}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\Psi}$ (9)
as in Proposition 1.7. We denote temporarily by $p_{A}\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$ also by $p_{A^{\prime}}\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$) the induced projection map w.r.t to $A$ (resp. to $A^{\prime}$).
  1. 1.
    Let $x\in M$. It is easily checked for all $V_{x}\in p_{A}^{-1}(x)$ and $V^{\prime}_{x}\in p_{A^{\prime}}^{-1}(x)$ that $\dim[{V}_{x}]=\dim[{V}_{x}^{\prime}]$, see [Lou23].
  2. 2.
    Let $x\in M$ and $k=\dim\mathrm{Sker}\rho_{x}$. Let $(u_{i})_{i=1,\ldots,k}$ be local sections of $A$ in an open neighborhood $\mathcal{U}_{x}$ of $x$ such that $\rho(u_{i})=0$ and $\mathrm{{Span}}(u_{i}(x))=\mathrm{Sker}\rho_{x}$. We can shrink $\mathcal{U}_{x}$ such that the map $y\mapsto\mathrm{Span}(u_{i}(y))\subseteq\mathrm{Sker}\rho_{y}$ be of constant rank on $\mathcal{U}_{x}$. Therefore, $K_{A}(x):=\sqcup_{y\in\mathcal{U}_{x}}\mathrm{Span}(u_{i}(y))\subset\ker\rho|_{\mathcal{U}_{x}}$ is a vector sub-bundle of $A$ of rank $k$ on $\mathcal{U}_{x}$ whose fiber over $x$ is $\mathrm{Sker}\rho_{x}$. Likewise, there exists a vector subbundle $K_{A^{\prime}}(x)$ of $A^{\prime}$ on $\mathcal{U}_{x}$ whose fiber over $x$ is $\mathrm{Sker}\rho^{\prime}_{x}$. Without any loss of generality, one can assume that $\Phi_{|_{y}}(K_{A}(x))\subseteq K_{A^{\prime}}(x)$ for all $y\in\mathcal{U}_{x}$. Thus, the map $\Phi_{|_{y}}$ goes to quotient
    $\phi_{y}\colon\frac{\ker\rho_{y}}{K_{A}(x)}\longrightarrow\frac{\ker\rho^{\prime}_{y}}{K_{A^{\prime}}(x)}$
    for all $y\in\mathcal{U}_{x}$ so that $\phi_{x}=[\Phi_{|_{x}}]$. Let $(x_{n})_{n\in\mathbb{N}}$ be a sequence of regular points converging to $x$ such that the sequence $\ker\rho_{x_{n}}$ converges to $V_{x}$ in $\mathrm{Grass}_{-r}(A)$. This implies that the sequence $\ker\rho_{x_{n}}/K_{A}(x)$ converges to $[V_{x}]=V_{x}/K_{A}(x)=V_{x}/\mathrm{Sker}\rho_{x}$ in $\mathrm{Grass}_{-r}\left(\frac{A}{K_{A}(x)}\right)$. Since $\phi_{x_{n}}\left(\frac{\ker\rho_{x_{n}}}{K_{A}(x)}\right)\subseteq\frac{\ker\rho^{\prime}_{x_{n}}}{K_{A^{\prime}}(x)}$, it follows that $\phi_{x}([V_{x}])\subseteq[V^{\prime}_{x}]$ where $V^{\prime}_{x}$ is the limit of (a sub-sequence of) $\ker\rho^{\prime}_{x_{n}}$. By item 1, $[V_{x}]$ and $[V^{\prime}_{x}]$ have the same dimension, therefore $\phi_{x}([V_{x}])=[V^{\prime}_{x}]$. This defines the required map and completes the proof.
2.2.1 Non-smoothness of $\mathrm{Nbl}_{\mathcal{F}}(M)$ and Nagano-Sussman
Although the blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)\subseteq\mathrm{Grass}_{-r}(A)$ may fail to be a smooth manifold (see Example 3.10), we explain how it can inherit a sort of “smooth structure”, i.e., it has a good enough framework to define the notion of singular foliations on $\mathrm{Nbl}_{\mathcal{F}}(M)$ that admits smooth leaves.
Since our blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ is defined in particular as a closed subset of $\mathrm{Grass}_{-r}(A)$, it can be equipped with a good notion of “smooth” functions
$C^{\infty}\left(\mathrm{Nbl}_{\mathcal{F}}(M)\right):=\left\{g_{|\mathrm{Nbl}_{\mathcal{F}}(M)},\,\text{for all}\,\,g\in C^{\infty}\left(\mathrm{Grass}_{-r}{(A)}\right)\right\}.$ (10)
If $\mathrm{Nbl}_{\mathcal{F}}(M)$ is a smooth submanifold, then it is an embedded submanifold since it is a closed subset. Moreover, the algebra $C^{\infty}\left(\mathrm{Nbl}_{\mathcal{F}}(M)\right)$ defined in (10) coincides with the algebra of smooth functions on $\mathrm{Nbl}_{\mathcal{F}}(M)$.
We introduce the following definitions, which are particular cases of a more general notion known as subcartesian differential spaces, see [Ś03]. We will work in the context of smooth differential geometry, but this can be adapted easily in the context of holomorphic/analytic cases.
  1. 1.
    Let $S$ be a closed subset of a manifold $N$. A vector field on $S\subseteq N$ is the restriction to $S$ of a vector fields $Z\in\mathfrak{X}(N)$ whose flow preserves $S$, i.e., $\phi^{Z}_{t}(S)\subseteq S$ whenever it makes sense. In that case, we shall say that such a $Z$ is tangent to $S$. The set of vector fields on $S$ form a Lie algebra that we denote as in the usual case by $\mathfrak{X}(S)$.
  2. 2.
    The tangent space $T_{s}S$ of $S$ at $s\in S$ is the evaluation at $s$ of the vector fields on $S$.
Remark 2.16.
Notice that when $S$ is a submanifold this notion of vector fields on $S$ agrees to the usual case.
Now, we recall a crucial theorem that allows to define singular foliations correctly on a closed subset $S$ of a manifold $N$.
Definition 2.17.
[LGLR24, LGLS20] Let $S$ be a closed subset of $N$.
  1. 1.
    A singular foliation on $S$ is an involutive [^4] Notice that $\mathcal{F}$ is generated by the restrictions to $S$ of vector fields on $N$, they are required to be involutive only after restrictions to $S$. locally finitely generated $\mathcal{O}_{S}$-submodule $\mathcal{F}\subseteq\mathfrak{X}(S)$.
  2. 2.
    For $s\in S$, the leaf of a singular foliation $\mathcal{F}$ on $S$ through $s$ is the set
    $L_{s}:=\left\{\phi_{t_{1}}^{Z_{1}}\circ\phi_{t_{2}}^{Z_{2}}\circ\cdots\circ\phi_{t_{k}}^{Z_{k}}(s),\;t_{1},\ldots,t_{k}\in\mathbb{R}\;\right\}$ (11)
    Above, $Z_{1},\ldots,Z_{k}$ are vector fields on $N$ whose restrictions are in $\mathcal{F}$. We implicitly assume that the flows are defined.
Remark 2.18.
From Definition 2.17, it is easily checked that being in the same leaf is an equivalence relation on $S$, hence the leaves induce a partition of $S$.
The notion of leaves of singular foliation on $S\subseteq M$ is justified by the following theorem that generalizes the Stefan-Sussman theorem [Ste74, Ste80], that says the leaves are smooth manifolds.
Theorem 2.19.
Let $\mathcal{F}$ be a singular foliation on a closed subset $S\subseteq M$. The leaves $\mathcal{F}$ form a partition of $S$ into connected manifolds, immersed as submanifolds of $M$.
The explanation of this result is based on a very strong theorem known as the Nagano-Sussman theorem [Nag66]. Let us first recall this theorem.
The following theorem, widely used in control theory, provides a very strong result regarding the smoothness of the orbits of a finite number of vector fields on a manifold without any assumptions.
Theorem 2.20 (Nagano–Sussmann).
Let $\mathcal{V}\subseteq\mathfrak{X}(N)$ be a locally finitely generated $\mathcal{O}_{N}$-submodule of vector fields on a manifold $N$. For every $\ell\in N$, the set
$\left\{\phi_{t_{1}}^{Z_{1}}\circ\phi_{t_{2}}^{Z_{2}}\circ\cdots\circ\phi_{t_{n}}^{Z_{n}}(\ell),\;t_{1},\ldots,t_{n}\in\mathbb{R},\;Z_{1},\ldots,Z_{n}\in\mathcal{V},\;n\in\mathbb{N}\right\}$
is a connected immersed submanifold of $N$.
Proof (of Theorem 2.19).
We can assume that $\mathcal{F}\subseteq\mathfrak{X}(S)$ is globally finitely generated. Let $\xi_{1},\ldots,\xi_{k}$ be generators for $\mathcal{F}$. By definition, the $\xi_{i}$’s are the restrictions to $S$ of vector fields $Z_{i}$’s on $N$ whose flows $\phi^{Z_{i}}_{t}$ preserves $S$, i.e., $\phi^{Z_{i}}_{t}(S)\subseteq S$ where the flows are defined. By Nagano-Sussman theorem, the orbits generated by the vector fields $Z_{1},\ldots,Z_{k}\in\mathfrak{X}(N)$ are immersed submanifolds of $N$. By assumption, the orbit that contains a point of $S$ is included in $S$. This completes the proof. ∎
Likewise, using this framework, the notion (almost) Lie algebroids can be naturally defined on a closed subset $S$ of a manifold $N$.
Definition 2.21.
An almost Lie algebroid over $S$ is the triplet $(A|_{S},[\cdot\,,\cdot]_{A,S},\rho_{A,S})$ where
  1. 1.
    $A|_{S}$ is the restriction of a vector bundle $A\to N$ to $S$;
  2. 2.
    $\rho_{A,S}$ is the restriction to $S$ of a vector bundle morphism $\rho_{A}\colon A\to TN$ such that $\rho(a)|_{S}$ defines a vector field on $S$, for all $a\in\Gamma(A)$;
  3. 3.
    $[\cdot\,,\cdot]_{A,S}$ is skew-symmetric $\mathbb{R}$-bilinear map on $\Gamma(A|_{S})$ obtained as the restriction of a skew-symmetric $\mathbb{K}$-bilinear map $[\cdot\,,\cdot]_{A}\colon\Gamma(A)\times\Gamma(A)\to\Gamma(A)$ satisfying for all $a,b\in\Gamma(A)$
    1. (a)
      the usual Leibniz identity w.r.t $\rho_{A}$ : $\left[a,fb\right]_{A}=\rho_{A}(a)[f]b+f\left[a,b\right]_{A}$,
    2. (b)
      $\rho_{A}([a,b]_{A})-[\rho_{A}(a),\rho_{A}(b)]\in\mathcal{I}_{S}\mathfrak{X}(N)$,
    where $\Gamma(A|S)$ are the restrictions of sections of $A$ to $S$ and $\mathcal{I}_{S}\subset\mathcal{O}_{N}$ is the ideal of vanishing functions on $S$.
If $[[a,b]_{A},c]_{A}+[[c,a]_{A},b]_{A}+[[b,c]_{A},a]_{A}\in\mathcal{I}_{S}\Gamma(A)$, for all $a,b,c\in\Gamma(A)$, we say $(A|_{S},[\cdot\,,\cdot]_{A,S},\rho_{A,S})$ is a Lie algebroid over $S$.
Remark 2.22.
For an almost Lie algebroid on a closed subset $S$ as in Definition 2.21, the image $\mathcal{F}=\rho_{A,S}(\Gamma(A|S))$ is a singular foliation on $S$ in the sense of Definition 2.17.
Example 2.23.
In this context, a Debord Lie algebroid on a closed subset $S\subseteq N$ is a Lie algebroid $(A|_{S},[\cdot\,,\cdot]_{A,S},\rho_{A,S})$ such that $\rho_{A,S}\colon A|_{S}\to TS$ is injective on an open dense subset of $S$. Also, A singular foliation $\mathcal{F}\subseteq\mathfrak{X}(S)$ is said to be a Debord foliation if and only if it is the image of a Debord Lie algebroid on $S$. In that case, $\mathcal{F}=\rho_{A,S}(\Gamma(A|_{S}))$ is a projective $\mathcal{O}_{S}$-module.
2.3 The (almost) Lie algebroid structure $\mathrm{Nash}(A)\to\mathrm{Nbl}_{\mathcal{F}}(M)$
In this section, we state the main results of the paper and discuss their proofs.
Recall that in Section 2.2, we constructed the blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ out of a finitely generated singular subalgebroid $\mathcal{F}\subseteq\Gamma(E)$, with the mapping
$p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\longrightarrow M,$
satisfying the conditions outlined in Proposition and 2.9 and Proposition 2.13 through an $E$-valued almost Lie algebroid $A\to M$, yet independent of it. In this section, we study the pullback of $A\to M$ to $\mathrm{Nbl}_{\mathcal{F}}(M)$. All constructions remain applicable for $E$-valued almost Lie algebroids. Subsequently, we narrow our focus to the case of Lie algebroids.
Notation 2.24.
In this section,
  1. 1.
    $(E,\left[\cdot\,,\cdot\right]_{E},\rho_{E})$ is a Lie algebroid over $M$,
  2. 2.
    $(A,\left[\cdot\,,\cdot\right]_{A},\rho)$ stands for an (almost) Lie algebroid valued in $E$, and its basic singular subalgebroid is denoted by $\mathcal{F}=\rho(\Gamma(A)))\subseteq\Gamma(E)$.
Recall that the regular part of $\mathcal{F}$ has constant dimension $r$. Also, $p^{!}A$ stands for the restriction to $\mathrm{Nbl}_{\mathcal{F}}(M)$ of the vector bundle $P^{!}A\to\mathrm{Grass}_{-r}(A)$. The pullback of a section $a\in\Gamma(A)$ to $\Gamma(p^{!}A)$ shall be denoted by $p^{!}a$.
2.3.1 Main statements
The main result of this paper establishes that any singular subalgebroid $\mathcal{F}\subseteq\Gamma(E)$ arising from a $E$-valued almost Lie algebroid $A\to M$ is desingularized into a Debord Lie algebroid over the blow-up space $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\to M$. This Lie algebroid embeds into an exact sequence, wherein the last two terms are almost Lie algebroids. The latter becomes a sequence of Lie algebroids if $A\to M$ is a Lie algebroid whose last term is Lie algebroid ideal.
More precisely,
Theorem 2.25.
The pull-back vector bundle $p^{!}A$ comes equipped with a “natural” almost Lie algebroid inscribed in an exact sequence
[…]
where $\mathcal{D}_{\mathcal{F}}$ is a Lie algebroid such that $\mathcal{D}_{\mathcal{F}}|_{p^{-1}(M_{\mathrm{reg}})}\simeq\mathcal{F}|_{M_{\mathrm{reg}}}$, and $K$ is a vector bundle whose fibers stable under the pointwise bracket of $p^{!}A$ and coincide with $\{\ker\rho_{m}\}_{m\in M}$ on regular points. By “natural” we mean that the projection $p^{!}A\to A$ is an almost Lie algebroid morphism. Also, $\mathcal{D}_{\mathcal{F}}\to\mathrm{Nbl}_{\mathcal{F}}(M)$ only depends on $\mathcal{F}$ not on $A$.
While establishing the proof of Theorem 2.25, we observe the following outcome concerning Lie algebroids and singular foliation, for the case where $E=TM$.
The next statement gives a precise meaning to the slogan “A Lie algebroid splits in a purely-manifold part and a purely Lie algebra part”. Recall that a Debord algebroid can be thought as a “purely-manifold” Lie algebroid, in the sense that it is entirely described by its image through the anchor map.
Theorem 2.26.
Let $E=TM$. For any Lie algebroid $A\to M$, the pull-back vector bundle $p^{!}A$ comes equipped with a “natural” Lie algebroid inscribed in an exact sequence
[…]
where $\mathcal{D}_{\mathcal{F}}$ is a Debord Lie algebroid such that $\mathcal{D}_{\mathcal{F}}|_{p^{-1}(M_{\mathrm{reg}})}\simeq\mathcal{F}|_{M_{\mathrm{reg}}}$, and $K$ is a Lie algebra bundle. By “natural” we mean that the projection $p^{!}A\to A$ is a Lie algebroid morphism. Also, $\mathcal{D}_{\mathcal{F}}\to\mathrm{Nbl}_{\mathcal{F}}(M)$ only depends on $\mathcal{F}$ not on $A$.
This makes sense of the following definition.
Definition 2.27.
The Nash blow-up $\mathrm{Nash}(A)\to\mathrm{Nbl}_{\mathcal{F}}(M)$ of an $E$-valued almost Lie algebroid/Lie algebroid $A\to M$ is the almost Lie algebroid/Lie algebroid structure on $p^{!}A$ of Theorem 2.25.
  1. 1.
    $\mathcal{D}_{\mathcal{F}}$ is the Lie algebroid/Debord Lie algebroid associated to $\mathcal{F}=\rho(\Gamma(A))\subseteq\Gamma(E)$;
  2. 2.
    $K$ is the tautological vector bundle/Lie algebra bundle associated to $A$.
The proof of Theorem 2.25 and 2.26 go through the assertions below.
Let us construct $K$ and $\mathcal{D}_{\mathcal{F}}$. The short exact sequence of vector bundles (7) on $\mathrm{Grass}_{-r}(A)$ restricts to a short exact sequence of vector bundles,
[…]
where $K$ and by $\mathcal{D}_{\mathcal{F}}$ stand for the restriction of the tautological bundle $\tau^{-r}_{A}$ (of codimension $r$) and quotient bundle $Q_{A}^{r}$ (of rank $r$) to $\mathrm{Nbl}_{\mathcal{F}}(M)$, respectively.
Proposition 2.28.
In the notations above:
  1. 1.
    For every $a\in\Gamma(A)$, there exists a unique vector field  $a^{\dagger}$ on $\mathrm{Nbl}_{\mathcal{F}}(M)$ such that $p_{*}(a^{\dagger})=\rho(a)\circ p$.
  2. 2.
    The mapping $a\in\Gamma(A)\mapsto a^{\dagger}\in\mathfrak{X}(\mathrm{Nbl}_{\mathcal{F}}(M))$ is a Lie algebroid action of $A$ on $\mathrm{Nbl}_{\mathcal{F}}(M)$, i.e., it satisfies for all $a,b\in\Gamma(A)$ and $f\in\mathcal{O}_{M}$
    • i.
      $[a,b]^{\dagger}=[a^{\dagger},b^{\dagger}]$,
    • ii.
      ${(fa)^{\dagger}}=(f\circ p)\,a^{\dagger}$,
    • iii.
      $p_{*}(a^{\dagger})=\rho(a)\circ p$.
  3. 3.
    there is a unique Lie algebroid structure on $p^{!}A$ such that the projection
    ${p^{!}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${A\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Nbl}_{\mathcal{F}}(M)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{p}$${M}$ (13)
    is a Lie algebroid morphism and whose anchor map is given by
    $\hat{\rho}(p^{!}a)=a^{\dagger},\;\;\text{for all $a\in\Gamma(A)$}.$
  4. 4.
    $K\rightarrow\mathrm{Nbl}_{\mathcal{F}}(M)$ is a Lie algebra bundle and $K\subset p^{!}A$ is a Lie algebroid ideal whose fibers coincide with the fibers of $\ker\rho$ on $p^{-1}(M_{\mathrm{reg}})$.
The following statement justifies the notation $\mathcal{D}_{\mathcal{F}}$.
Corollary 2.29.
In the notations of Proposition 2.28, there is a Lie algebroid structure $\left(\mathcal{D}_{\mathcal{F}},\,[\cdot,\cdot\,]_{\mathcal{D}_{\mathcal{F}}},\,\overline{\rho}\right)$ on the quotient bundle $\mathcal{D}_{\mathcal{F}}$. If $E=TM$, its anchor map $\overline{\rho}\colon\mathcal{D}_{\mathcal{F}}\to T\left(\mathrm{Nbl}_{\mathcal{F}}(M)\right)$ is injective on the open dense subset $p^{-1}(M_{\mathrm{reg}})$, i.e., it is a Debord algebroid.
Proposition 2.28 and Corollary 2.29 implies that
$0\rightarrow K\rightarrow p^{!}A\rightarrow\mathcal{D}_{\mathcal{F}}\rightarrow 0$ (14)
is a short exact sequence of Lie algebroids. This proves Theorem 2.26.
Remark 2.30.
For $E=TM$, the blowup map $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\longrightarrow M$ sends the leaves of $\mathcal{D}_{\mathcal{F}}$ to the leaves of $\mathcal{F}$ and the restriction $p|_{p^{-1}(M_{\mathrm{reg}})}$ is an isomorphism of regular foliations. In particular, the Nash blow-up does not change the dimension of regular leaves.
We end this section with a description of leaves of $\mathcal{D}_{\mathcal{F}}$ with $E=TM$. Let $L\subseteq M$ be a leaf of the Lie algebroid $A\to M$. The restriction $A|_{L}\to TL$ is a transitive Lie algebroid over $L$. The latter acts on $p^{-1}(L)\hookrightarrow\mathrm{Grass}_{-(r-\dim L)}(\ker\rho|_{L})$. The orbits of this action are those of $\mathcal{D}_{\mathcal{F}}|_{p^{-1}(L)}$ by construction. Let us describe these leaves.
For simplicity, we assume that $A$ is integrable to a source-simply connected Lie groupoid $\Gamma\rightrightarrows M$. For $L\subseteq M$, we shall denote $\Gamma^{L}_{L}:=s^{-1}(L)\cap t^{-1}(L)$ and $\Gamma^{m}_{m}:=\Gamma^{\{m\}}_{\{m\}}$, where $s,t$ are source and target of $\Gamma$.
Proposition 2.31.
Let $p^{!}\mathcal{F}$ be the basic singular foliation of $\mathcal{D}_{\mathcal{F}}$ and $L\subset M$ a leaf of $\mathcal{F}$. The restriction groupoid $\Gamma^{L}_{L}\rightrightarrows L$ to $L$ naturally acts on $\mathrm{Grass}_{-(r-\dim L)}(\ker\rho|_{L})$.
  1. 1.
    this action preserves the fiber
    $p^{-1}(L)\hookrightarrow\mathrm{Grass}_{-(r-\dim L)}(\ker\rho|_{L})$
    of the blow-up map $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$.
  2. 2.
    If $L=\{m\}$ is a leaf of $\mathcal{F}$, then the leaf $L_{V_{m}}$ of $p^{!}\mathcal{F}$ that passes through $V_{m}\in p^{-1}(m)$ corresponds to a $\Gamma^{L}_{L}$-orbit and we have $L_{V_{m}}\simeq\frac{\Gamma^{m}_{m}}{\mathrm{Stab}(V_{m})}$. Moreover,
    1. (a)
      $V_{m}\in p^{-1}(m)$ is a leaf of $p^{!}\mathcal{F}$ if only if $V_{m}$ is a Lie ideal of $\mathfrak{g}_{m}(\mathcal{F})$.
    2. (b)
      $V_{m}\in p^{-1}(m)$ is a regular point of $p^{!}\mathcal{F}$ if only if $\mathrm{Stab}(V_{m})=V_{m}$.
    where $\mathrm{Stab}(V_{m})$ is the stabilizer of $V_{m}\in p^{-1}(m)$. In that case, the Debord Lie algebroid $\mathcal{D}_{\mathcal{F}}|_{p^{-1}(m)}$ is
    $\dfrac{\Gamma^{m}_{m}\times\mathfrak{g}_{m}(\mathcal{F})/{V_{m}}}{N(V_{m})}\to\frac{\Gamma^{m}_{m}}{N(V_{m})}$
    where $N(V_{m})$ is the normalizator of with $V_{m}\in p^{-1}(m)$.
2.3.2 Proofs of main results
In this section, we prove Proposition 2.28 and Corollary 2.29. The latter are proved through Proposition 1.8, Lemma 2.32 and Lemma 2.33.
Here, we write the proofs in the context $E=TM$, i.e., for almost Lie algebroids over singular foliations $\mathcal{F}\subseteq\mathfrak{X}(M)$. The singular subalgebroid case is similar.
Lemma 2.32 (Lift of vector fields of $\mathcal{F}$ on $\mathrm{Nbl}_{\mathcal{F}}(M)$).
In the notations of Proposition 1.8, consider a lift $\hat{\xi}_{X}$ of some $X\in\mathcal{F}$ to the Grassmann bundle $\mathrm{Grass}_{-r}(A)$ of some anchored bundle $(A,\rho)$, where $\xi_{X}$ is a linear vector field on $A$ over $X$
  1. 1.
    the flow $\phi_{t}^{\hat{\xi}_{X}}$ of $\hat{\xi}_{X}$, whenever it is defined, preserves the subset $\mathrm{Nbl}_{\mathcal{F}}(M)\subset\mathrm{Grass}_{-r}(A)$, i.e.
[…]
  2. 2.
    In particular, $\hat{\xi}_{X}$ is tangent to $\mathrm{Nbl}_{\mathcal{F}}(M)$ is the neighborhood of any point where $\mathrm{Nbl}_{\mathcal{F}}(M)$ is a sub-manifold.
Proof.
For every $x\in M$, the flow $\phi_{t}^{\xi_{X}}$ (whenever it makes sense) sends $\mathrm{\ker}{\rho}_{x}$ to $\ker{\rho}_{\phi^{X}_{t}(x)}$, that is
$\phi_{t}^{\xi_{X}}|_{x}\left(\mathrm{\ker}{\rho}_{x}\right)=\ker{\rho}_{\phi^{X}_{t}(x)}$
in view of Diagram (4). In particular, in terms of the section $\sigma_{\ker}:M_{\mathrm{reg}}\to\mathrm{Grass}_{-r}(A)$,  $x\mapsto\ker\rho_{x}$, it means that
$\phi^{\hat{\xi}_{X}}_{t}\circ\sigma_{\ker}=\sigma_{\ker}\circ\phi^{X}_{t}.$
In particular, since $\phi^{X}_{t}(M_{\mathrm{reg}})\subseteq M_{\mathrm{reg}}$, this implies that $\phi^{\hat{\xi}_{X}}_{t}$ preserves the closure $\overline{\sigma_{\ker}(M_{\mathrm{reg}})}$, i.e., it preserves $\mathrm{Nbl}_{\mathcal{F}}(M)$. This completes the proof. ∎
Lemma 2.33.
In the smooth case, there exists a vector bundle morphism
${\tilde{\rho}\colon P^{!}A\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${T(\mathrm{Grass}_{-r}(A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Grass}_{-r}(A)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathrm{Grass}_{-r}(A)}$
such that
  1. 1.
    The following diagram commutes:
[…]
  2. 2.
    For any $a\in\Gamma(A)$, $\tilde{\rho}(P^{!}a)\in\mathfrak{X}\left(\mathrm{Grass}_{-r}(A)\right)$ is a lift of $X=\rho(a)\in\mathcal{F}$ tangent to $\mathrm{Nbl}_{\mathcal{F}}(M)$. Here $P^{!}a$ is the pull-back of the section $a$.
Here $P^{!}A$ stands for the pull-back of $A$ on the manifold $\mathrm{Grass}_{-r}(A))$ and $P^{!}A\rightarrow A$ is the projection onto $A$. In the holomorphic/real analytic cases, existence is local.
Proof.
Let $V_{x}\in{\mathrm{Grass}}_{-r}(A)$ be an element in the fiber of $x\in M$. Consider $a_{1},\dots,a_{d}$ a local trivialization of $A$ in a neighborhood $\mathcal{U}$ of $x$. The evaluation at $V_{x}\in{\mathrm{Grass}}_{-r}(A)$ of the lifts $\hat{\xi}_{a_{1}},\dots,\hat{\xi}_{a_{i}}:=\hat{\xi}_{{\rho}(a_{i})},\cdots\hat{\xi}_{a_{d}}$ constructed as in Lemma 2.32 define a vector bundle morphisms $\tilde{\rho}_{\mathcal{U}}$ which satisfies both required properties above by construction. In the smooth case, using a partition of unity $(\mathcal{U}_{i},\phi_{i},\rho_{\mathcal{U}_{i}})_{i\in I}$, this construction can be made global: it is easily checked that $\sum_{i}P^{*}\phi_{i}\tilde{\rho}_{\mathcal{U}_{i}}$ still satisfies the required properties 1) and 2) because each $\tilde{\rho}_{\mathcal{U}_{i}}$ defined as before does. ∎
Proof of Proposition 2.28 and 2.31
Proof of Proposition 2.28.
  1. 1.
    Let $a\in\Gamma(A)$ and consider $\rho(a)=X$. In the notations of Lemma 2.32, let $\hat{\xi}_{X}$ be the lift of $X$ to $\mathrm{Grass}_{-r}(A)$. The restriction of $\hat{\xi}_{X}$ to ${\mathrm{Nbl}}_{\mathcal{F}}(M)$ is the required vector field. Uniqueness is a direct consequence of the fact that $p$ is an isomorphism when restricted to the dense open subset $p^{-1}(M_{\mathrm{reg}})$.
  2. 2.
    The construction of the map $a\mapsto a^{\dagger}$ is obtained by restricting the vector bundle morphism
    $P^{!}A\stackrel{{\scriptstyle{\rho}}}{{\rightarrow}}T\left(\mathrm{Grass}_{-r}(A)\right)$
    of Lemma 2.33 to $\mathrm{Nbl}_{\mathcal{F}}(M)$. The vector fields $[a,b]_{A}^{\dagger}$ and $[a^{\dagger},b^{\dagger}]$ are both $p$-related to $[\rho(a),\rho(b)]$, for all $a,b\in\Gamma(A)$. By uniqueness, we have
    $[a,b]_{A}^{\dagger}=[a^{\dagger},b^{\dagger}],\,\,\hbox{for all $a,b\in\Gamma(A)$}.$
  3. 3.
    The Lie algebroid structure on $p^{!}A$ is obtained as follows: the anchor map $\hat{\rho}$ is defined by $p^{!}a\mapsto a^{\dagger}$ and the Lie bracket on $p^{!}A$ as
    $[p^{!}a,p^{!}b]_{p!A}:=p^{!}([a,b]_{A})$
    for all $a,b\in\Gamma(A)$. We extend by Leibniz identity to all sections.
  4. 4.
    The fact that $K\rightarrow\mathrm{Nbl}_{\mathcal{F}}(M)$ is a Lie algebra bundle is a direct consequence of item 3 of Proposition 2.9. Let us show that it is a Lie algebroid ideal of $p^{!}A$. Notice that $\Gamma(K)\subset\ker\hat{\rho}$ and $\Gamma(K)|_{p^{-1}(M_{\mathrm{reg}})}=\ker\hat{\rho}|_{p^{-1}(M_{\mathrm{reg}})}$ : indeed, we have $\ker\hat{\rho}_{V_{x}}\subseteq\ker\rho_{x}$ for all $x\in M$ by construction, and both spaces coincide on $p^{-1}(M_{\mathrm{reg}})$ since $\dim(\mathrm{Im}\tilde{\rho}_{V_{x}})=\dim(\mathrm{Im}{\rho}_{x})$ for all $x\in M_{\mathrm{reg}}$. For $E\neq TM$, we only have the inclusion $\Gamma(K)\subset\ker\hat{\rho}$.
    Now, take $\xi\in\Gamma(K)\subset\ker\hat{\rho}$,   $\ell\in\Gamma(p^{!}A)$ and $V\in\mathrm{Nbl}_{\mathcal{F}}(M)$. One has,
    $\hat{\rho}\left([\xi,\ell]_{p^{!}A}\right)=[\hat{\rho}(\xi),\hat{\rho}(\ell)]=0.$
    Let $V_{n}\in p^{-1}(M_{\mathrm{reg}})$ be a sequence that converges to $V$. For all $n\in\mathbb{N}$, we have
    $[\xi,\ell]_{p^{!}A}(V_{n})\in\ker\hat{\rho}_{V_{n}}=\ker{\rho}_{V_{n}}=V_{n}.$
    This implies that the limit $[\xi,\ell]_{p^{!}A}(V_{n})\underset{n\to+\infty}{\longrightarrow}[\xi,\ell]_{p^{!}A}(V)$ belongs to $V$. This proves that
    $\left[\Gamma(K),\;\Gamma(p^{!}A)\right]_{p^{!}A}\subseteq\Gamma(K).$
Proof of Corollary 2.29.
We use notations of Proposition 2.28. By Proposition 2.28(4), $K\subset p^{!}A$ is a Lie algebroid ideal, therefore the anchor map $\hat{\rho}\colon p^{!}A\mapsto T\left(\mathrm{Nbl}_{\mathcal{F}}(M)\right)$ goes to the quotient
[…]
to equip $\mathcal{D}_{\mathcal{F}}$ with a Lie algebroid structure. If $E=TM$, the fibers of the tautological bundle $K\to\mathrm{Nbl}_{\mathcal{F}}(M)$ coincide with the fibers $\{\ker\hat{\rho}_{V_{x}}\}_{V_{x}\in p^{-1}(x),x\in M}$ on the open dense subset $p^{-1}(M_{\mathrm{reg}})$, $\mathcal{D}_{\mathcal{F}}$ is a Debord Lie algebroid.∎
Proof (of Proposition 2.31).
We give a proof with a leaf $L=\{m\}$ reduced to a point, the general case is similar. Let $(A,[\cdot\,,\cdot]_{A},\rho)$ be a Lie algebroid and $m\in M$. Notice that for every $a\in\Gamma(A)$, the flow $\phi^{\xi_{X}}_{t}\colon A\rightarrow A$ of the linear vector field $\xi_{X}\in\mathfrak{X}(A)$ associated to $\mathcal{L}_{a}$, with $\rho(a)=X$, satisfies for all $b\in\Gamma(A)$
$\frac{d}{dt}_{|_{t=0}}(\phi^{\xi_{X}}_{t})^{*}b=[a,b]_{A}$ (17)
with
$(\phi^{\xi_{X}}_{t})^{*}(b)(m):=\phi^{\xi_{X}}_{-t}(m)(b(\phi^{X}_{t}(m))),$
whenever the flows make sense. Using the formula (17) we make the following observation: let $G_{m}(\mathcal{F})$ be the simply connected Lie group integrating $\mathfrak{g}_{m}(\mathcal{F})$. The latter naturally acts on $\mathrm{Grass}_{-(r-\dim(L_{m})}\left(\mathfrak{g}_{m}(\mathcal{F})\right)$ through the adjoint action. By Lemma 2.32 (1), this action preserves the fiber
$p^{-1}(m)\hookrightarrow\mathrm{Grass}_{-(r-\dim(L_{m})}\left(\mathfrak{g}_{m}(\mathcal{F})\right)$
of the blow-up map $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$. If $m\in M$ is a leaf of $\mathcal{F}$, then the leaf $L_{V_{m}}$ of $p^{!}\mathcal{F}$ that passes through $V_{m}\in p^{-1}(m)$ is a ${G}_{m}(\mathcal{F})$-orbit, since the flows $\phi^{\xi}_{t}$ generates the foliation $p^{!}\mathcal{F}$. It is easily checked that
  1. 1.
    $V_{m}\in p^{-1}(m)$ is a leaf of $p^{!}\mathcal{F}$ if only if $[V_{m}]$ is a Lie ideal of $\mathfrak{g}_{m}(\mathcal{F})$.
  2. 2.
    Since
    $\dim L_{V_{m}}=r-\frac{\mathrm{Stab}([V_{m}])}{[V_{m}]},$
    $V_{m}\in p^{-1}(m)$ is a regular point if only if $\mathrm{Stab}([V_{m}])=[V_{m}]$.
3 Examples
In this section, we present some examples related to singular foliations, Lie algebroids, and also to Poisson manifolds. The main results of the paper in Section 2 provide a desingularization method for $E$-valued (almost) Lie algebroids over a manifold $M$. They can also be used to desingularize finitely generated singular subalgebroids $\mathcal{F}\subseteq\Gamma(E)$, particularly finitely generated singular foliations $\mathcal{F}\subseteq\mathfrak{X}(M)$ by choosing an almost Lie algebroid $A$ over $\mathcal{F}$. The blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ and the Debord foliation $\mathcal{D}_{\mathcal{F}}$ as in Theorem 2.25 are independent of the choice of the almost Lie algebroid by Proposition 2.15. Our primary focus lies in computing the blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ and the Debord algebroid $p\colon\mathcal{D}_{\mathcal{F}}\to M$ associated with $\mathcal{F}$. We shall denote the basic singular foliation of $\mathcal{D}_{\mathcal{F}}$ by $p^{!}\mathcal{F}$.
3.1 Lie algebroids, Lie algebra actions, singular foliations
Example 3.1 (The Nash-blow-up does not affect projectivity).
If $(A,[\cdot\,,\cdot]_{A},\rho)$ is a Lie algebroid whose anchor map is injective on an open dense subset, then
[…]
since $\mathrm{Grass}_{-\mathrm{rk}(A)}(A)\simeq M$.
For instance, the Nash blow-up has no effect on the Lie algebroid of vector fields on a $b$-manifold in the sense of [GMP14].
Example 3.2 (Vector fields tangent to a closed submanifold).
Let $M$ be a manifold and $W\subseteq M$ a closed submanifold. We denote by $\mathcal{F}_{W}=\mathfrak{X}_{W}(M)$ the singular foliation of all vector fields tangent to $W$ [LGLR24]. It is still an open question whether $\mathcal{F}_{W}$ is the image through the anchor map of a Lie algebroid on the whole manifold $M$. However, the answer is positive in a tubular neighborhood of $W$ in $M$. Indeed, such neighborhood comes equipped with a diffeomorphism to a neighborhood of the zero section of the normal bundle $\mathcal{N}_{W}=\frac{TM|_{W}}{TW}$. Under this diffeomorphism, the restriction of $\mathcal{F}_{W}$ to that neighborhood corresponds to the space of all vector fields on $\mathcal{N}_{W}$ tangent to the zero section, that we denote again by $\mathcal{F}_{W}$. It follows that:
  1. 1.
    Linear vector fields on $\mathcal{N}_{W}$ are the sections of a transitive Lie algebroid over $W$ known as $\mathrm{CDO}(\mathcal{N}_{W})$ in [Mac05],
  2. 2.
    $\mathrm{CDO}(\mathcal{N}_{W})$ acts on $\mathcal{N}_{W}$,
  3. 3.
    $\mathcal{F}_{W}$ is the image the anchor map of this action Lie algebroid on $\mathcal{N}_{W}$.
For simplicity, we can assume that the tubular neighborhood is $M$. It is easy to check that $p\colon\mathrm{Nbl}_{\mathcal{F}_{W}}(M)\rightarrow M$ is the usual blow-up of $M$ along $W$ (see, e.g., [Huy04]) and $p^{-1}(W)\simeq\mathbb{P}(\mathcal{N}_{W})$. The pullback $p^{!}\mathcal{F}_{W}\simeq\mathcal{D}_{\mathcal{F}_{W}}$ of $\mathcal{F}$ along $p\colon\mathrm{Nbl}_{\mathcal{F}_{W}}(M)\rightarrow M$ is the Debord singular foliation made of all vector fields on $\mathrm{Nbl}_{\mathcal{F}_{W}}(M)$ that are tangent to $p^{-1}(W)$.
In conclusion, we have the short exact sequence of Lie algebroids over the blowup of $\mathbb{K}^{d}$ at 00 as follows
$0\to K\to p^{!}(\mathrm{CDO}(\mathcal{N}_{W}))\to\mathcal{D}_{\mathcal{F}_{W}}\to 0,$
here $K$ is made of pairs $(\ell,\mathfrak{M})\in\mathbb{P}(\mathcal{N}_{W})\times\mathfrak{gl}(\mathcal{N}_{W})$ such that $\ell$ is a an eigenspace of $\mathfrak{M}$.
The following example is a particular case of Example 3.2 with $W$ being a single point.
Example 3.3 (Action of $\mathfrak{gl}_{d}(\mathbb{K})$ on $M=\mathbb{K}^{d}$).
Consider the action Lie algebroid $A=\mathfrak{gl}_{d}(\mathbb{K})\ltimes\mathbb{K}^{d}\to\mathbb{K}^{d}$. Its basic singular foliation $\mathcal{F}=\mathrm{Span}(X_{kj}=x_{k}\frac{\partial}{\partial x_{j}},\;k,j=1,\ldots,d)$ is generated by of all vector fields vanishing at the origin. It is easily checked that the Nash blow-up $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\to\mathbb{K}^{d}$ coincides to the usual blow-up of $\mathbb{K}^{d}$ at the origin. In the $i$-th coordinate chart $(\mathcal{U}_{i},y_{1}\ldots,y_{d})$ of the blow-up space, the pullback $p^{!}\mathcal{F}$ is generated by the $\frac{\partial}{\partial y_{j}}$’s with $j\neq i$ and $y_{i}\frac{\partial}{\partial y_{i}}.$ These computations show that $p^{!}\mathcal{F}\simeq\mathcal{D}_{\mathcal{F}}$ is generated by all vector fields tangent to the projective space $p^{-1}(0)=\mathbb{P}^{d-1}$.
We have the short exact sequence of Lie algebroids
$0\to K\to p^{!}(\mathfrak{gl}_{d}(\mathbb{K})\ltimes\mathbb{K}^{d})\to\mathcal{D}_{\mathcal{F}}\to 0,$
here $K$ is made of pairs $(\ell,\mathfrak{M})\in\mathbb{P}^{d-1}\times\mathfrak{gl}_{d}(\mathbb{K})$ such that $\ell$ is a an eigenspace of $\mathfrak{M}$.
Example 3.4 (Vector fields vanishing at order $k\in\mathbb{N}$ on $M=\mathbb{K}^{d}$).
Denote by $x_{1},\ldots,x_{d}$ the canonical coordinates of $\mathbb{K}^{d}$. Let $k\in\mathbb{N}$ and $\mathcal{F}^{k}$ be the singular foliation on $\mathbb{K}^{d}$ generated by the vector fields $x^{I}\frac{\partial}{\partial x_{j}}$ for $j=1,\ldots,d$ and $x^{I}$ ranges over all monomials of degree $k$ in $x_{1},\ldots,x_{d}$. For every $k\in\mathbb{N}$, $M_{\mathrm{reg}}=\mathbb{K}^{d}\setminus\{0\}$. The singular foliation $\mathcal{F}^{k}$ is the image of an almost Lie algebroid $A=\mathbb{K}^{d}\times\mathbb{K}^{N}$ whose anchor map $\rho$ is $e_{I,j}\mapsto x^{I}\frac{\partial}{\partial x_{j}}$ and the almost Lie bracket is given by
$[e_{I,j},e_{J,l}]_{A}:=x^{J\setminus j}e_{I,l}-x^{I\setminus l}e_{J,j}$ (18)
with $N=d\times\begin{pmatrix}d+k\\ k\end{pmatrix}$ and $x^{I\setminus j}$ is the monomial $\frac{x^{I}}{x_{j}}$. For $k=1$, $\mathcal{F}^{1}$ is given by the action $\mathfrak{gl}_{d}(\mathbb{K})\curvearrowright\mathbb{K}^{d}$, see Example 3.3. For $k\neq 1$, it is an open question whether $\mathcal{F}^{k}$ is the image of some Lie algebroid or not, see [LGLS20].
A direct computation shows that for every $k\in\mathbb{N}$,
$\mathrm{Nbl}_{\mathcal{F}^{k}}(M)\simeq\mathrm{Bl}_{\mathcal{I}_{k}}(\mathbb{K}^{d}).$
where $\mathrm{Bl}_{\mathcal{I}_{k}}(\mathbb{K}^{d})$ is the usual blow-up of $\mathbb{K}^{d}$ along the ideal $\mathcal{I}_{k}$ of all polynomial functions vanishing at order $k$. Consider the embedding
$\nu_{k}\colon\mathbb{P}^{d-1}\rightarrow\mathbb{P}^{N-1},\,[x_{1}:\cdots:x_{d}]\mapsto[\cdots:x^{I}:\cdots]$ (19)
known as the Veronese map of degree $k$, see e.g., [Har92]. For $k=2$ and $d=2$, the map (19) reads
$\nu_{2}\colon\mathbb{P}^{1}\rightarrow\mathbb{P}^{2},\,[x:y]\mapsto[x^{2}:xy:y^{2}].$
Under the embedding (19), $\mathrm{Bl}_{\mathcal{I}_{k}}(\mathbb{K}^{d})$ is isomorphic to the usual blow-up of $\mathbb{K}^{d}$ at 00. Also, $p_{k}^{-1}(0)$ is isomorphic to $\nu_{k}(\mathbb{P}^{d-1})$, where $p_{k}\colon\mathrm{Nbl}_{\mathcal{F}^{k}}(M)\rightarrow M$ is the projection map.
Example 3.5 ($\mathfrak{sl}_{2}\curvearrowright\mathbb{R}^{2}$).
Consider the Lie algebra $\mathfrak{sl}_{2}\subset\mathfrak{gl}_{2}$ generated by the canonical generators $h,e,f$ that satisfy $[h,e]=2e$, $[h,f]=-2f$ and $[e,f]=h$. It acts on $\mathbb{R}^{2}$ as: $\rho\colon\mathbb{R}^{2}\times\mathfrak{sl}_{2}\rightarrow T\mathbb{R}^{2}$
$\rho\begin{pmatrix}h\\ e\\ f\end{pmatrix}(x,y)=\begin{pmatrix}x&-y\\ 0&x\\ y&0\end{pmatrix}\begin{pmatrix}\frac{\partial}{\partial x}\\ \frac{\partial}{\partial y}\end{pmatrix}$
Here, $x,y$ stand for the coordinates of $M=\mathbb{R}^{2}$. The set of regular points is $M_{\mathrm{reg}}=\mathbb{R}^{2}\setminus\{0\}$. The blow-up space $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\subset\mathbb{R}^{2}\times\mathrm{Grass}_{-2}(\mathfrak{sl}_{2})\to\mathbb{R}^{2}$ is the usual blow-up of $\mathbb{R}^{2}$ along the ideal $I=(-x^{2},xy,y^{2})$. This corresponds to the usual blowup $\mathrm{Bl}_{0}(\mathbb{R}^{2})$ of $\mathbb{R}^{2}$ at 00: indeed, we have $\nu\left(\mathrm{Bl}_{0}(\mathbb{R}^{2})\right)=\mathrm{Nbl}_{\mathcal{F}}(M)$ under the closed embedding
$\nu\colon\mathbb{R}^{2}\setminus\{0\}\times\mathbb{P}^{1}\hookrightarrow\mathbb{R}^{2}\setminus\{0\}\times\mathbb{P}^{2},\;\left((x,y),[x:y]\right)\mapsto\left((x,y),[-x^{2}:xy:y^{2}]\right).$ (20)
We also know the blow-up space is smooth by applying directly Proposition 2.11: the kernel $\ker\rho$ is a projective $\mathcal{O}_{\mathbb{R}^{2}}$-submodule: It is the line bundle generated by the section $xyh+y^{2}e-x^{2}f$, see Example 3.31 of [LGLS20]. By a direct computation one can find the generators of the Debord singular foliation $p^{!}\mathcal{F}$, e.g., on the $x$-chart $\mathcal{U}_{x}$ of the blow-up space, it is generated by the vector fields $\overline{h}=x\frac{\partial}{\partial x}-2y\frac{\partial}{\partial y}$,  $\overline{e}=\frac{\partial}{\partial y}$ and $\overline{f}=xy\frac{\partial}{\partial x}-y^{2}\frac{\partial}{\partial y}$. Since $\overline{f}=y\overline{h}+y^{2}\overline{e}$, the Debord singular foliation $p^{!}\mathcal{F}$ is generated by the $\mathcal{O}_{\mathbb{R}^{2}}$-linearly independent vector fields $\overline{h},\,\overline{e}$ on $\mathcal{U}_{x}$. Here, $p^{-1}(0)=\{x=0\}$ is the unique singular leaf of $p^{!}\mathcal{F}$.
3.2 Foliated bivector fields
In this section, we give examples of our blow-up construction in Poisson geometry.
For a bivector field $\pi=\{\cdot\,,\cdot\}\in\mathfrak{X}^{2}(M)$ we denote by $\pi^{\#}\colon T^{*}M\longrightarrow TM$ the vector bundle morphism which is defined by the relation $\langle\pi^{\#}(\alpha),\beta\rangle=\langle\pi,\alpha\wedge\beta\rangle$ with $\alpha,\beta\in\Omega^{1}(M)$.
The result of this section can be applied to a more general concept than Poisson structures. Let us recall the following definition introduced in [Tur15].
Definition 3.6.
A bivector field $\pi\in\mathfrak{X}^{2}(M)$ is said to be foliated if the submodule $\pi^{\#}(\Omega^{1}(M))\subset\mathfrak{X}(M)$ is a singular foliation, i.e.,
$\left[\pi^{\#}(\Omega^{1}(M)),\pi^{\#}(\Omega^{1}(M))\right]\subset\pi^{\#}(\Omega^{1}(M)).$ (21)
The singular foliation $\pi^{\#}(\Omega^{1}(M))$ shall be denoted by $\mathcal{F}_{\pi}$ and be called the basic singular foliation of $\pi$.
Example 3.7.
Poisson structures on manifolds [LGPV13, CFM21] are the first class of examples of foliated bivector fields. Recall that a Poisson structure on a manifold $M$ is a bivector field $\pi=\{\cdot\,,\cdot\}\in\mathfrak{X}^{2}(M)$ fulfilling $[\pi,\pi]_{\mathrm{NS}}=0$ where $[\cdot\,,\cdot]_{\mathrm{NS}}$ is the Schouten–Nijenhuis bracket on multivector fields. For every function $h\in\mathcal{O}_{M}$ we associated a vector field $X_{h}:=\pi^{\#}(dh)=\{\cdot\,,h\}\in\mathfrak{X}(M)$ called the Hamiltonian vector field of $h$. One has that $[X_{h},X_{g}]=X_{\{g,h\}}$ for all $g,h\in\mathcal{O}_{M}$. This implies that every Poisson structure $\pi$ on a manifold is foliated. Another class of examples of foliated bivector fields are twisted-Poisson structures, i.e.,$[\pi,\pi]_{\mathrm{NS}}=\wedge^{3}\pi^{\#}(\alpha)$, where $\alpha\in\Omega^{3}(M)$ is a closed 3-form [vW01, Tur15].
For any foliated bivector field $(M,\pi)$, $\pi^{\#}\colon T^{*}M\rightarrow TM$ is anchored bundle over $\mathcal{F}_{\pi}$ and thus can be equipped with an almost Lie bracket structure $[\cdot\,,\cdot]_{\pi}$. From now on, we consider $\pi^{\#}\colon T^{*}M\rightarrow TM$ with an almost Lie algebroid structure over $\mathcal{F}_{\pi}$. When $\pi$ is Poisson [CFM21] or twisted Poisson[vW01], in both cases there exists Lie algebroid structure on cotangent bundle $T^{*}M$ whose anchor map is $\pi^{\#}\colon T^{*}M\rightarrow TM$. Therefore, Theorem 2.26 is applicable in all these cases.
In the notations of Section 2.2, the following proposition shows that for every foliated bivector field $(M,\pi)$ the Nash construction $\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}_{\pi}}(M)\subset\mathrm{Grass}_{r}(TM)$ out of the image and the one $\mathrm{Nbl}^{\ker}_{\mathcal{F}_{\pi}}(M)\subset\mathrm{Grass}_{-r}(T^{*}M)$ out of the kernel are canonically isomorphic.
Proposition 3.8.
Let $(M,\pi)$ be a (foliated) bivector field and $\pi^{\#}\colon T^{*}M\rightarrow TM$ its associated almost Lie algebroid over $\mathcal{F}_{\pi}$. Then $\mathrm{Nbl}^{\ker}_{\mathcal{F}_{\pi}}(M)\simeq\mathrm{Nbl}^{\mathrm{Im}}_{\mathcal{F}_{\pi}}(M)$.
Proof.
There is a canonical bijection between $\mathrm{Grass}_{-(d-r)}(TM)$ and $\mathrm{Grass}_{-r}(T^{*}M)$ which maps, $V\subset T_{x}M$ to its annihilator $V^{\circ}\subset T^{*}_{x}M$, with $d=\dim M$. On the other hand, since $\pi$ is skew-symmetric, i.e., $0=\langle\pi^{\#}(\alpha),\beta\rangle=-\langle\alpha,\pi^{\#}(\beta)\rangle$ for all $\alpha\in\ker(\pi^{\#}_{x})$ and $\beta\in T^{*}_{x}M$, the annihilator of $\mathrm{Im}(\pi^{\#}_{x})$ is $\ker(\pi^{\#}_{x})$
Example 3.9 (Concentric spheres).
Consider the Lie–Poisson structure on $M=~{}{\mathfrak{so}}(3,\mathbb{R})^{*}$ given on a dual basis $x,y,z$ by
$\pi=-z\frac{\partial}{\partial x}\wedge\frac{\partial}{\partial y}+y\frac{\partial}{\partial x}\wedge\frac{\partial}{\partial z}-x\frac{\partial}{\partial y}\wedge\frac{\partial}{\partial z}.$
The singular foliation $\mathcal{F}_{\pi}$ is generated by the vector fields
$X:=\pi^{\#}(dx)=z\frac{\partial}{\partial y}-y\frac{\partial}{\partial z},\;\;Y:=\pi^{\#}(dy)=z\frac{\partial}{\partial x}-x\frac{\partial}{\partial z},\;\;Z:=\pi^{\#}(dz)=y\frac{\partial}{\partial x}-x\frac{\partial}{\partial y}.$
The open dense subset of regular points is $M_{\mathrm{reg}}=\mathbb{R}^{3}\setminus\left\{0\right\}$. The kernel of $\pi^{\#}$ at a point $(x,y,z)\in M_{\mathrm{reg}}$ is the line generated by $(x,y,z)$. Hence, the blow-up space $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)=\overline{\left\{(m,\ell)\in\mathbb{R}^{3}\setminus\left\{0\right\}\times\mathbb{P}^{2}\,|\,m\in\ell\right\}}$ is the usual blow-up $\mathrm{Bl}_{0}(\mathbb{R}^{3})$ of $\mathbb{R}^{3}$ at 00. In particular, $\mathrm{Nbl}_{\mathcal{F}}(M)$ is a smooth manifold.
Let us look at the blow-up space for example in the $x$-chart given by $(x,y,z)\mapsto\left((x,xy,xz),[1:y,z]\right)$ so that $p((x,xy,xz),[1:y,z])=(x,xy,xz)$. The singular foliation $p^{!}\mathcal{F}_{\pi}$ is generated by the vector fields
$\overline{X}=z\frac{\partial}{\partial y}-y\frac{\partial}{\partial z},\quad\overline{Y}=xz\frac{\partial}{\partial x}-yz\frac{\partial}{\partial y}-(z^{2}+1)\frac{\partial}{\partial z}$
$\text{and}\;\;\overline{Z}=xy\frac{\partial}{\partial x}-(y^{2}+1)\frac{\partial}{\partial y}-yz\frac{\partial}{\partial z}$. Since $\overline{X}=y\overline{Y}-z\overline{Z}$, it follows that $p^{!}\mathcal{F}_{\pi}$ is a regular foliation on $\mathrm{Nbl}_{\mathcal{F}}(M)$ of rank $2$. We have a short exact sequence of Lie algebroids
$0\rightarrow O(-1)_{\mathrm{Nbl}_{\mathcal{F}}(M)}\rightarrow p^{!}\left(\mathfrak{so}(3,\mathbb{R})\ltimes\mathbb{R}^{3}\right)\rightarrow p^{!}\left(\mathfrak{so}(3,\mathbb{R})\ltimes\mathbb{R}^{3}\right)/O(-1)_{\mathrm{Nbl}_{\mathcal{F}}(M)}\rightarrow 0$
where $O(-1)_{\mathrm{Nbl}_{\mathcal{F}}(M)}$ denotes the Lie algebra bundle obtained by the restriction of the tautological line bundle $O(-1)\rightarrow\mathbb{P}^{2}\times\mathbb{R}^{3}$ to $\mathrm{Nbl}_{\mathcal{F}}(M)$ and
$p^{!}\mathcal{F}_{\pi}\simeq p^{!}\left(\mathfrak{so}(3,\mathbb{R})\ltimes\mathbb{R}^{3}\right)/O(-1)_{\mathrm{Nbl}_{\mathcal{F}}(M)}.$
The bivector field $\pi$ does not lift to $\mathrm{Nbl}_{\mathcal{F}}(M)$. Recall that the blowup space of $\mathbb{R}^{3}$ at the origin $\mathrm{Bl}_{0}(\mathbb{R}^{3})\subset~{}\mathbb{P}^{2}\times\mathbb{R}^{3}$ is covered by three charts given by $x\neq 0$, $y\neq 0$ and $z\neq 0$. Let us look at the $x$-chart where the projection $p$ becomes $(x,y,z)\mapsto~{}(x,xy,xz)$. In this chart, $\pi$ pulls back to
$y\frac{\partial}{\partial z}\wedge\frac{\partial}{\partial x}+z\frac{\partial}{\partial x}\wedge\frac{\partial}{\partial y}+\frac{1}{x}(1+y^{2}+z^{2})\frac{\partial}{\partial y}\wedge\frac{\partial}{\partial x}.$ (22)
For $x=0$, Equation (22) is not defined. In conclusion, the Hamiltonian vector fields of the Poisson structure $(M,\pi)$ lift to $\mathrm{Nbl}_{\mathcal{F}}(M)$ which is a smooth manifold, but the bivector field $\pi$ does not lift to a smooth bivector field on $\mathrm{Nbl}_{\mathcal{F}}(M)$.
In fact, the Poisson structure of Example 3.9 can be seen as a particular case of a more general construction of Poisson structures in dimension $3$.
Example 3.10.
Let $\varphi\in\mathbb{K}[x,y,z]$. Consider the Poisson structure on $\mathbb{K}^{3}$ given by:
$\pi=\frac{\partial\varphi}{\partial z}\frac{\partial}{\partial x}\wedge\frac{\partial}{\partial y}-\frac{\partial\varphi}{\partial y}\frac{\partial}{\partial x}\wedge\frac{\partial}{\partial z}+\frac{\partial\varphi}{\partial x}\frac{\partial}{\partial y}\wedge\frac{\partial}{\partial z}$
This Poisson structure was studied in [Pic06]. Consider $\mathcal{F}_{\pi}$ its the singular foliation generated by the vector fields $X_{ij}=\frac{\partial\varphi}{\partial x_{i}}\frac{\partial}{\partial x_{j}}-\frac{\partial\varphi}{\partial x_{j}}\frac{\partial}{\partial x_{i}}$ with $x_{i},x_{j}\in\{x,y,z\}$. When $\frac{\partial\varphi}{\partial x}$, $\frac{\partial\varphi}{\partial y}$ and $\frac{\partial\varphi}{\partial z}$ form a complete intersection, $\mathcal{F}_{\pi}$ is exactly the space of all $X\in\mathfrak{X}(\mathbb{K}^{3})$ satisfying $X[\varphi]=0$, see [LGLS20] Section 3.3. The set of the singular points is the zero locus $W$ of the ideal generated by $\frac{\partial\varphi}{\partial x}$, $\frac{\partial\varphi}{\partial y}$ and $\frac{\partial\varphi}{\partial z}$. For $m=(x,y,z)\in M_{\mathrm{reg}}$, a direct computation shows that $\ker\pi^{\#}|_{m}=\left[\frac{\partial\varphi}{\partial x}(m):\frac{\partial\varphi}{\partial y}(m):\frac{\partial\varphi}{\partial z}(m)\right]$. Therefore, $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)$ is the natural transformation of $\mathbb{K}^{3}$ along $W$ which is not smooth in general. Take, e.g., $\varphi=xy-\frac{z^{n+1}}{n+1}$ for some integer $n\geq 2$: if we denote by $[u:v:w]$ the coordinates of the projective space $\mathbb{P}^{2}$, $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)=\{uy-vx=0,wx-uz^{n}=0,wy-vz^{n}=0\}$ is the blow-up of $\mathbb{K}^{3}$ along the ideal $(x,y,z^{n})$. In the chart $u=1$, $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)=A^{n}:wx-z^{n}=0$ is a Du Val singularity.
By Theorem 2.26, even when $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)$ is singular, it admits a Debord singular foliation $p^{!}\mathcal{F}_{\pi}$ whose leaves are manifolds included in $\mathrm{Nbl}_{\mathcal{F}_{\pi}}(M)$.
We end this section with an example where the blow-up space $\mathrm{Nbl}_{\mathcal{F}}(M)$ arises as a quotient space.
Example 3.11 (The adjoint action of $\mathfrak{su}(n)$).
Consider the adjoint action of the Lie algebra $\mathfrak{su}(n)$ of traceless anti-Hermitian $n\times n$ complex matrices. These matrices are diagonalizable in a unitary basis. $M_{\mathrm{reg}}\subset\mathfrak{su}(n)$ is the open dense subset of matrices whose centralizers have minimal dimension, i.e., traceless anti-Hermitian matrices that admit distinct eigenvalues. By definition $\mathrm{Nbl}_{\mathcal{F}}(M)=\overline{\{C(\mathfrak{M}),\mathfrak{M}\in M_{\mathrm{reg}}\subset\mathfrak{su}(n)\}}$, here $C(\mathfrak{M})$ is the centralizer of $\mathfrak{M}\in\mathfrak{su}(n)$.
Denote by $\mathfrak{D}\subseteq\mathfrak{su}(n)$ the space of diagonal anti-Hermitian matrices. We have a well-defined surjective map
$\frac{{SU}(n)\times\mathfrak{D}}{\mathrm{Stab}{(\mathfrak{D}})}\to\mathfrak{su}(n),\;[(U,D)]\mapsto U^{-1}DU,$
which is proper and invertible on $M_{\mathrm{reg}}$. Notice that for all $\mathfrak{M}\in M_{\mathrm{reg}}$ such that $\mathfrak{M}=U^{-1}DU$ for some $(U,D)\in{SU}(n)\times\mathfrak{D}$, we have $C(U^{-1}DU)=U^{-1}\mathfrak{D}U$. This implies that there is a well-defined map at the level of the Grassmanian
$\frac{{SU}(n)\times\mathfrak{D}}{\mathrm{Stab}{(\mathfrak{D}})}\to\mathrm{Nbl}_{\mathcal{F}}(M),\;[(U,D)]\mapsto(U^{-1}DU,\,U^{-1}\mathfrak{D}U).$ (23)
Here, $SU(n)$ stands for the Lie group of $n\times n$ unitary matrices with determinant 1, and $\mathrm{Stab}{(\mathfrak{D}})=\{U\in{SU}(n)\mid U^{-1}\mathfrak{D}U=\mathfrak{D}\}$ is the stabilizator of $\mathfrak{D}$ in $\mathfrak{su}(n)$. The result is similar for $\mathfrak{so}(n)$. Let us show that the map (23) is a bijection:
  1. 1.
    Surjectivity: let $V\in p^{-1}(\mathfrak{M})$ and a convergent sequence of regular points $\mathfrak{M}_{n}\underset{n\to+\infty}{\longrightarrow}\mathfrak{M}$ such that $C(\mathfrak{M}_{n})\underset{n\to+\infty}{\longrightarrow}V$. We have that $\mathfrak{M}_{n}=U_{n}^{-1}D_{n}U_{n}\underset{n\to+\infty}{\longrightarrow}~{}\mathfrak{M}$ with $(U_{n},D_{n})\in{SU}(n)\times\mathfrak{D}$. By compacity of $SU(n)$, there exists a subsequence of $U_{n}^{-1}\mathfrak{D}U_{n}$ that converges to $V$. This proves surjectivity.
  2. 2.
    Injectivity: let $U_{1},U_{2}\in SU(n)$ be such that $U_{1}^{-1}\mathfrak{D}U_{1}=U_{2}^{-1}\mathfrak{D}U_{2}$. This implies that $(U_{1}U_{2}^{-1})^{-1}\mathfrak{D}(U_{1}U_{2}^{-1})=\mathfrak{D}$. Thus, $U_{1}U_{2}^{-1}\in\mathrm{Stab}{(\mathfrak{D}})$.
A more general result than Example 3.11 is the topic of a forth-coming paper.
4 Conclusion
We end the paper with a natural question. This paper highlights for a given Lie algebroid $(A,[\cdot,\,\cdot]_{A},\rho_{A})$ a procedure for separating the space of objects (i.e., the singular foliation part) and the symmetry part (a Lie algebra bundle). That is, $A$ admits a pullback $p^{!}A$ along the blow-up space $p\colon\mathrm{Nbl}_{\mathcal{F}}(M)\rightarrow M$ that comes with a short exact sequence of Lie algebroids
$0\rightarrow K\rightarrow p^{!}A\rightarrow\mathcal{D}_{\mathcal{F}}\rightarrow 0$ (24)
with $K$ a totally intransitive Lie algebroid (i.e., a Lie algebra bundle) and $\mathcal{D}_{\mathcal{F}}$ a Lie algebroid whose anchor map is injective on an open dense subset.
It is natural to ask whether $p^{!}A$ is a semi-direct product $\mathcal{D}_{\mathcal{F}}\ltimes K$, in other words:
Question 4.1.
What is the meaning of the curvature $C\colon\wedge^{2}\mathcal{D}_{\mathcal{F}}\rightarrow K$,
$C(u,v):=s([u,v]_{\mathcal{D}_{\mathcal{F}}})-[s(u),s(v)]_{p^{!}A}$
of a section $s\colon\mathcal{D}_{\mathcal{F}}\rightarrow p^{!}A$ of $p^{!}A\rightarrow\mathcal{D}_{\mathcal{F}}$
Notice that any section $s\colon\mathcal{D}_{\mathcal{F}}\rightarrow p^{!}A$ of (24) gives rise to a Lie algebra morphism
$\mathcal{D}_{\mathcal{F}}\rightarrow\mathrm{Der}(Z(K)),\;\alpha\mapsto\nabla_{\alpha}\beta:=[s(\alpha),\beta]_{p^{!}A}$
i.e., a representation of $\mathcal{D}_{\mathcal{F}}$ on the center $Z(K)$ of $K$ [ABR18]. Moreover, if the isotropy Lie algebras $\{\ker\rho_{x}\}_{x\in M}$ of $A$ are Abelian, then $K$ is Abelian and $(\mathcal{D}_{\mathcal{F}},\nabla)$ is a representation of $\mathcal{D}_{\mathcal{F}}$ on $K$.